### v1

# Analysis of Corrected Graph Convolutions

Abstract
Machine learning for node classification on graphs is a prominent area driven by applications such as recommendation systems. State-of-the-art models often use multiple graph convolutions on the data, as empirical evidence suggests they can enhance performance. However, it has been shown empirically and theoretically, that too many graph convolutions can degrade performance significantly, a phenomenon known as oversmoothing. In this paper, we provide a rigorous theoretical analysis, based on the contextual stochastic block model (CSBM), of the performance of vanilla graph convolution from which we remove the principal eigenvector to avoid oversmoothing. We perform a spectral analysis for $k$ rounds of corrected graph convolutions, and we provide results for partial and exact classification. For partial classification, we show that each round of convolution can reduce the misclassification error exponentially up to a saturation level, after which performance does not worsen. For exact classification, we show that the separability threshold can be improved exponentially up to $O({\log{n}}/{\log\log{n}})$ corrected convolutions.
1 Introduction
Graphs naturally represent complex relational information found in a plethora of applications such as social analysis  (Backstrom and Leskovec, 2011), recommendation systems (Ying et al., 2018; Borisyuk et al., 2024), computer vision (Monti et al., 2017), materials science and chemistry (Reiser et al., 2022), statistical physics (Battaglia et al., 2016; Bapst et al., 2020), financial forensics (Zhang et al., 2017; Weber et al., 2019) and traffic prediction in Google Maps (Derrow-Pinion et al., 2021).
The abundance of relational information in combination with features of the corresponding entities has led to improved performance of machine learning models for classification and regression tasks. Central to the field of machine learning on graphs is the graph convolution operation. It has been shown empirically (Defferrard et al., 2016; Kipf and Welling, 2017; Gasteiger et al., 2019; Rossi et al., 2020) that using graph convolutions to the feature data enhances the prediction performance of a model, but too many graph convolutions can have the opposite effect (Oono and Suzuki, 2020; Chen et al., 2020; Keriven, 2022; Wu et al., 2023), an issue known as oversmoothing. Several solutions have been proposed for this problem, we refer the reader to the survey of Rusch et al. (2023).
In this paper we provide a rigorous spectral analysis, based on the contextual stochastic block model (Deshpande et al., 2018), to show that the oversmoothing phenomenon can be alleviated by excluding the principal eigenvector’s component from the graph convolution matrix. This is similar to the state-of-the-art normalization approach used in Zhao and Akoglu (2020); Rusch et al. (2023) (PairNorm), however, our method explicitly uses the principal eigenvector. We provide below some intuition about why excluding the principal eigenvector helps to alleviate over-smoothing.
Let $A$ be the adjacency matrix of the given graph, and $D$ be the degree matrix. Vanilla graph convolutions are represented using matrices such as $D^{-1}A$ or ${D^{-1/2}AD^{-1/2}}$ (Kipf and Welling, 2017). Suppose our graph is $d-$regular, meaning that each node has exactly $d$ neighbors. In this case, both graph convolutions reduce to $\frac{1}{d}A$. The top eigenvector of $A$ is ${\mathbbm{1}}$ with eigenvalue $d$, where ${\mathbbm{1}}$ is the vector of all ones. This means that $\lim_{k\rightarrow\mathcal{\infty}}\frac{1}{d^{k}}A^{k}=\frac{1}{n}{\mathbbm{1}}{\mathbbm{1}}^{\top}$, which implies that multiplying our feature vector, $x$, by $\frac{1}{d}A$ many times is essentially equivalent to projecting our data onto the all-ones vector, $x\mapsto\frac{1}{n}\langle x,{\mathbbm{1}}\rangle{\mathbbm{1}}$. This means that all feature values will converge to the same point. Therefore, we should expect, as verified by most real-world and synthetic experiments, that many rounds of the convolution $x\mapsto\frac{1}{d}Ax$ will lead to a large learning error. However, if we instead perform convolution with the corrected matrix $\tilde{A}:=\frac{1}{d}A-\frac{1}{n}{\mathbbm{1}}{\mathbbm{1}}^{\top}$, then the convergent behavior of $x\mapsto\tilde{A}^{k}x$ would be equivalent to projecting $x$ onto the second eigenvector of $A$. This eigenvector is known to capture information about sparse bipartitions in the graph $G$ (Cheeger, 1970; Alon and Milman, 1985; Alon, 1986), and so for certain problems, like binary classification, we may expect this eigenvector to capture a larger amount of information about our signal.
In our analysis, we will work on the problem of linear binary classification in the contextual stochastic block model, and our results are stated in terms of the following corrected convolution matrices:
$\hat{A}$ $=D^{-1/2}AD^{-1/2}-D^{1/2}\mathbbm{1}\mathbbm{1}^{\top}D^{1/2}\qquad\mbox{ and }\qquad\tilde{A}=\frac{1}{d}A-\frac{1}{n}{\mathbbm{1}}{\mathbbm{1}}^{\top},$ (1)
where $d$ is the empirical average degree in $A$, which is equal to $2|E|/n$ and $|E|$ is the number of edges in the graph. Note that $\hat{A}$ is derived from the normalized adjacency matrix, while $\tilde{A}$ is (up to a scalar multiple) its unnormalized counterpart. Briefly, in our results, we demonstrate that when the graph is sufficiently dense and of reasonably good quality, the corrected graph convolutions exponentially improve both partial and exact classification guarantees. Depending on the density and quality of the given graph, improvement becomes saturated after $\mathcal{O}(\log n)$ convolutions in our partial and exact classification results. However, in comparison to a similar analysis in (Wu et al., 2023) for vanilla graph convolutions (without correction), we show that classification accuracy does not become worse as the number of convolutions increases.
1.1 Our Contributions
In this work, we provide, to our knowledge, the first theoretical guarantees on partial and exact classification after $k$ rounds of graph convolutions in the contextual stochastic block model. Our main result is to show that each graph convolution with the corrected matrix reduces the classification error by a multiplicative factor until a certain point of “saturation” and the number of convolutions required until saturation depends on the amount of input feature variance. We show that the accuracy of the linear classifier at the point of saturation only depends on the strength of the signal from the graph. This is in contrast to the uncorrected convolution matrix, which will always exhibit a decrease in classification accuracy after many convolutions. Finally, we show that given slightly stronger assumptions on graph density and signal strength, the convolved data at the point of saturation will be linearly separable with high probability. To quantify our results, we let $p$ and $q$ be the intra- and inter-class edge probabilities with $\gamma(p,q)={|p-q|}/{(p+q)}$ being the “relative signal strength” in the graph. Let $\bar{d}$ be the expected degree of each vertex. Our results are as follows:
  • If $p+q\geq\Omega(\frac{\log^{2}{n}}{n})$ and $\gamma\geq\Omega(\frac{1}{\sqrt{\bar{d}}})$, each convolution with $\tilde{A}$ reduces classification error by a factor of about $\frac{1}{\gamma^{2}\bar{d}}$ until the fraction of errors is $O(\frac{1}{\gamma^{2}\bar{d}})$
  • If $p+q\geq\Omega(\frac{\log^{2}{n}}{n})$ and $\gamma\geq\Omega(\sqrt{\frac{\log{n}}{\bar{d}}})$, each convolution with $\hat{A}$ reduces classification error by a factor of about $\frac{\log{n}}{\gamma^{2}\bar{d}}$ until the fraction of errors is $O(\frac{\log{n}}{\gamma^{2}\bar{d}})$
  • If $p+q\geq\Omega(\frac{\log^{3}{n}}{n})$, $\gamma\geq\Omega(k\sqrt{\frac{\log{n}}{\bar{d}}})$ and the input features has signal-to-noise ratio at least $\Omega(\sqrt{\frac{\log{n}}{n}})$, the data is linearly separable after $k$ rounds of convolutions with $\tilde{A}$.
To obtain our partial classification results, we use spectral analysis to bound the mean-squared-error between the convolved features and the true signal. For exact classification, we prove a concentration inequality on the total amount of message received by a vertex through “incorrect paths” of length $k$ after $k$ rounds of convolution through a combinatorial moment analysis. Using this, we establish entry-wise bounds on the deviation of the convolved feature vector from the true signal.
2 Literature review
Research on graph learning has increasingly focused on methods that integrate node features with relational information, particularly within the semi-supervised node classification framework, see, for example, Scarselli et al. (2009); Cheng et al. (2011); Gilbert et al. (2012); Dang and Viennet (2012); Günnemann et al. (2013); Yang et al. (2013); Hamilton et al. (2017); Jin et al. (2019); Mehta et al. (2019); Chien et al. (2022); Yan et al. (2021). These studies have underscored the empirical advantages of incorporating graph structures when available.
The literature also addresses the expressive capacity (Lu et al., 2017; Balcilar et al., 2021) and generalization potential (Maskey et al., 2022) of Graph Neural Networks (GNNs), including challenges like oversmoothing (Keriven, 2022; Xu et al., 2021; Oono and Suzuki, 2020; Li et al., 2018; Rusch et al., 2023). In our paper, we ground our work on the contextual stochastic block model (Deshpande et al., 2018), a widely used statistical framework for analyzing graph learning and inference problems. Recent theoretical studies have extensively used the CSBM to illustrate several statistical, information-theoretic, and combinatorial results on relational data accompanied by node features. In Deshpande et al. (2018); Lu and Sen (2020), the authors investigate the classification thresholds for accurately classifying a significant portion of nodes from this model, given linear sample complexity and large but bounded degree. Additionally, Hou et al. (2020) introduces graph smoothness metrics to quantify the utility of graphical information. Further developments in Chien et al. (2021, 2022); Baranwal et al. (2021, 2023a) extend the application of CSBM, establishing exact classification thresholds for graph convolutions in multi-layer networks, accompanied by generalization guarantees. A theoretical exploration of the graph attention mechanism (GAT) is provided by Fountoulakis et al. (2023), delineating conditions under which attention can improve node classification tasks. More recently, Baranwal et al. (2023b) provide the locally Bayes optimal message-passing architecture for node classification for the general CSBM.
In this paper, we provide exact and partial classification guarantees for multiple graph convolution operations. Previous investigations have often been confined to a few convolution layers, limiting the understanding of their effects on variance reduction (see, for example, Baranwal et al. (2023a)). Our findings contribute a novel spectral perspective on graph convolutions, describing how the fraction of recoverable nodes is influenced by the signal-to-noise ratio in the node features and the scaled difference between intra- and inter-class edge probabilities. We also demonstrate that the oversmoothing phenomenon can be alleviated by excluding the principal eigenvector’s component from the adjacency matrix – a strategy somewhat akin to the normalization approach used in Zhao and Akoglu (2020) (PairNorm), albeit our method explicitly uses the principal eigenvector and is grounded in rigorous spectral justifications.
In a relatively recent work (Wu et al., 2023) the authors rigorously analyze the phenomenon of oversmoothing in GNNs for the 2-block CSBM by identifying two competing effects of graph convolutions: the mixing effect, which homogenizes node representations across different classes, and the denoising effect, which homogenizes node representations within the same class. Their analysis shows that oversmoothing occurs when the mixing effect dominates the denoising effect, and they quantify the number of layers required for this transition. In contrast, we work with the corrected graph convolution in the 2-block CSBM and show that it improves performance exponentially up to saturation, after which more convolutions do not improve nor degrade performance. On a technical level, the previous work only analyzes the distribution of a single node’s feature values after convolution and does not take into account correlations between nodes. In our work, we use spectral analysis of higher powers of the convolution matrix, which takes into account correlations between nodes to obtain our partial classification results over the whole dataset. To handle the modified convolution in the exact classification setting, we analyze the error more directly through matrix perturbation analysis rather than trying to directly count the higher-order neighbors of each vertex as in previous works (Baranwal et al., 2023a; Wu et al., 2023).
3 Preliminaries and Model Description
Throughout this paper, we use $\mathbbm{1}$ to denote the all-ones vector and $e_{i}$ to denote the $i^{th}$ standard basis vector in $\mathbb{R}^{n}$. Given a vector $x\in\mathbb{R}^{n}$, we use $\left\lVert x\right\rVert$ to denote its Euclidean norm $\sqrt{\sum_{i=1}^{n}x(i)^{2}}$. We use $\left\lVert x\right\rVert_{\infty}$ to denote its infinity norm, $\max_{i=1}^{n}|x(i)|$. For a matrix $M\in\mathbb{R}^{n}$, we use $\left\lVert M\right\rVert$ to denote its spectral norm, $\max_{x\neq 0,\left\lVert x\right\rVert=1}\left\lVert Mx\right\rVert$. We also make routine use of the spectral theorem, which says that if $M$ is a $n\times n$ symmetric matrix, then it can be diagonalized with $n$ orthogonal eigenvectors and real eigenvalues. In particular, there exist $\lambda_{1},\lambda_{2},...\lambda_{n}\in\mathbb{R}$ and orthonormal vectors $w_{1},w_{2},...w_{n}\in\mathbb{R}^{n}$ such that $M=\sum_{i=1}^{n}\lambda_{i}w_{i}w_{i}^{\top}$. Note that when $M$ is symmetric, $\left\lVert M\right\rVert=\max_{i}|\lambda_{i}|=\max_{x:\left\lVert x\right\rVert=1}|x^{\top}Mx|$. Finally, we use the $\mathcal{N}(\mu,\Sigma)$ to a Gaussian distribution with mean $\mu$ and covariance matrix $\Sigma$. For one-dimensional Gaussians, we use $\mathcal{N}(\mu,\sigma^{2})$. For $X\sim\mathcal{N}(\mu,\sigma^{2})$, we will frequently use the Gaussian tail bound: $\mathbf{Pr}\left[|X-\mu|>t\sigma\right]\leq\exp(-\frac{t^{2}}{2})$.
3.1 Contextual Stochastic Block Model
In this section, we formally describe the contextual stochastic block model introduced by  (Deshpande et al., 2018). Our model is defined by parameters $n,m\in\mathbb{N}$, $p,q\in[0,1]$, $\mu,\nu,\in\mathbb{R}^{m}$ and $\sigma\in\mathbb{R}^{+}$ In the model, we are given a random undirected graph, $G=(V,E)$, where $|V|=n$, drawn from the 2-block stochastic block model and features drawn from the Gaussian mixture model. Our vertices are partitioned into two classes, $S$ and $T$, of size $n/2$, which we want to recover. For each pair of vertices $i,j\in S$ and $i,j\in T$, the edge $(i,j)$ is in $E$ independently with probability $p$ while for each pair $i\in S$ and $j\in T$, the edge $(i,j)$ is in $E$ with probability $q$. In addition to the graph, we are also given a feature matrix $X\in\mathbb{R}^{n\times m}$ drawn from a Gaussian mixture model with two centers $\mu$ and $\nu$. For each $i\in V$, we let $g_{i}\sim\mathcal{N}(0,\sigma^{2}I_{m})$ be an i.i.d. Gaussian noise vector. For $i\in S$, the $i^{th}$ row of $X$ is given by $\mu+g_{i}$, while for each $j\in T$, the $j^{th}$ row of $X$ is given by $\nu+g_{j}$.
4 Results and Interpretation
In our analysis, there are two types of objectives. In the exact classification objective, the aim is to exactly recover $S$ and $T$ with probability $1-o(1)$. In the partial classification, or “detection” problem, the goal is to correctly classify $1-o(1)$ fraction of vertices correctly with probability $1-o(1)$. We begin by stating our results for the partial classification regime. For ease of notation, we will assume $p>q$ from this point forward. We show in Appendix B that this assumption is made without loss of generality.
Theorem 4.1.
Suppose we are given a 2-block $m$-dimensional CSBM with parameters $n,p>q,\mu,\nu,\sigma$ satisfying $\gamma(p,q):=\frac{p-q}{p+q}\geq\Omega\left(\sqrt{\frac{1}{np}}\right)$ and $p\geq\Omega\left(\frac{\log^{2}{n}}{n}\right)$. There exists a linear classifier such that after $k$ rounds of convolution with $\tilde{A}$, will, with probability at least $1-\frac{1}{2}\exp(-\Omega\Big{(}\frac{n\left\lVert\mu-\nu\right\rVert^{2}}{\sigma^{2}}\Big{)})$, misclassify at most
$O\left(\frac{1}{\gamma^{2}p}+\Big{(}\frac{C}{\gamma\sqrt{np}}\Big{)}^{2k}\frac{\sigma^{2}}{\left\lVert\mu-\nu\right\rVert^{2}}n\log{n}\right)$
vertices, where $C$ is an absolute constant. Furthermore, if $\gamma\geq\Omega(\sqrt{\frac{\log{n}}{np}})$ then with probability at least $1-\frac{1}{2}\exp(-\Omega\Big{(}\frac{n\left\lVert\mu-\nu\right\rVert^{2}}{\sigma^{2}}\Big{)})$, the same linear classifier after $k$ rounds of convolution with $\hat{A}$ will misclassify at most
$O\left(\frac{\log{n}}{\gamma^{2}p}+\Big{(}\frac{C\log{n}}{\gamma\sqrt{np}}\Big{)}^{2k}\frac{\sigma^{2}}{\left\lVert\mu-\nu\right\rVert^{2}}n\log{n}\right)$
vertices.
Now we take a closer look at the error bound. For brevity, we will focus on our results regarding convolutions with $\tilde{A}$. First, we see that an important ratio in our bound is the term ${1}/({\gamma^{2}np})$. This term is small if $\gamma^{2}$ is much larger than the inverse of the expected degree of each vertex, $(p+q)n/2$, which is at most $np$. Our assumption that this ratio is upper bounded by a constant means that we need the signal from the graph to be sufficiently strong. Now, if we examine our misclassification error bound, and let $\rho={C}/({\gamma^{2}np})$ where $C$ is a sufficiently large constant, then we see that the fraction of misclassified vertices is at most $\rho+\rho^{k}\sigma^{2}\log{n}/\left\lVert\mu-\nu\right\rVert^{2}$. Our assumption on the parameters ensures that $\rho<1$. Note that only the second term depends on $k$, and the feature’s noise-to-signal ratio. This term measures the amount of error introduced by the variance in the features and exponentially decreases with $k$. Moreover, after about $k=\log_{1/\rho}\Big{(}{\sigma^{2}\log{n}}/({\rho\left\lVert\mu-\nu\right\rVert^{2}})\Big{)}$ convolutions, the $\rho$ term, which only depends on graph parameters, will dominate over the variance term, indicating that more convolutions will not improve the quality of the convolved features beyond the quality of the signal from the graph. If $\sigma/\left\lVert\mu-\nu\right\rVert$ is constant, we will always reach our optimal error bound of $O(\rho)$ when $k=O(\log\log{n})$. Moreover, if $\gamma=\Omega(1)$, as was assumed in Baranwal et al. (2021), then we will have $1/\rho\geq\Omega(np)$. This means even when $\sigma/\left\lVert\mu-\nu\right\rVert\approx\sqrt{n/\log{n}}$, we will reach optimality in constant number of convolutions with high probability if the graph is moderately dense. For example, if $p=1/\sqrt{n}$, then we only need $3$ convolutions and if $p=\Omega(1)$, then we only need $2$. On the other hand, if $\gamma$ is on the order of $\Theta(1/\sqrt{np})$, then in the worst case, we may need $\log{n}$ convolutions to reach our optimal bound. Next, we state our results for exact classification.
Theorem 4.2.
Suppose we are given a 2-block $m$-dimensional CSBM with parameters $n,p>q,\mu,\nu,\sigma$ satisfying $\gamma(p,q)\geq\Omega\left(k\sqrt{\frac{\log{n}}{np}}\right)$ and $p\geq\frac{\log^{3}{n}}{n}$. Then after $k=O(\log{n})$ rounds of graph convolution with $\tilde{A}$, our data is linearly separable with probability $1-n^{-\Omega(1)}$ if
$\frac{\left\lVert\mu-\nu\right\rVert}{\sigma}\geq\Omega\left(\max\left(\sqrt{\frac{\log{n}}{n}},\;\Big{(}\frac{C}{\gamma\sqrt{np}}\Big{)}^{k}\sqrt{\log{n}}\right)\right)$
where $C$ is an absolute constant.
Here, we bound the minimum signal-to-noise ratio required for exact classification as a function of $p,q,n$ and $k$. Just like the partial classification result, our function has a term that decreases exponentially with $k$ and a term independent of $k$. The rate of decrease for the dependent term is proportional to ${1}/({\gamma\sqrt{np}})$, or $\sqrt{\rho}$. We see once again that with more convolutions, the requirement on the feature signal-to-noise ratio for exact classification becomes exponentially weaker. Moreover, since we assumed that $\gamma\geq\Omega(k\sqrt{{\log{n}}/({np})})$, as long as $\left\lVert\mu-\nu\right\rVert\geq\Omega(\sigma\sqrt{{\log{n}}/{n}})$, the data becomes linearly separable after $k=O({\log{n}}/{\log\log{n}})$ convolutions. Just as in the partial classification case, we observe that the larger $\gamma$ is, the fewer convolutions we need to obtain the optimal bound. In particular, if $\gamma=\Omega(1)$ and $p=\Omega(1)$ then one convolution already gives the optimal bound, and if $p=1/\sqrt{n}$, then two convolutions are enough. For technical reasons, we only analyze exact classification using convolution with the corrected un-normalized adjacency matrix, $\tilde{A}$. Similar bounds should hold for $\hat{A}$ based on our simulation results, but we leave this for future work.
4.1 Discussion on our Assumptions
Both Theorem 4.1 and Theorem 4.2 require a lower bound of $\gamma\geq\omega({1}/{\sqrt{np}})$, and this is to ensure that the signal from the graph is strong enough so that a convolution does not destroy the signal from the data. Also, implicit in the probability bound of Theorem 4.1 and in Theorem 4.2, is the assumption that our signal-to-noise ratio, $\left\lVert\mu-\nu\right\rVert/\sigma$, is at least $\omega(1/\sqrt{n})$ so the feature noise does not completely drown out the signal. Our lower-bound assumption on $p$ is to ensure concentration in the behavior of the degrees and the adjacency matrix towards their expectation. In Theorem 4.2, we also assume that $k=O(\log{n})$. This is done mainly for technical reasons of our proof but we note that this assumption is made without loss of generality because as mentioned, the bound in Theorem 4.2 does not improve for $k\geq\log{n}$. Finally, we note that the case $p>q$ corresponds to a homophilous graph, and the case $p<q$ corresponds to a heterophilous graph (see Luan et al. (2021); Ma et al. (2022) for more). For binary classification, it has been shown (Baranwal et al., 2023a) that one can assume $p>q$ without loss of generality and make corresponding adjustments in the classifier. As such, we assume that $p>q$. For more detail regarding this assumption, see Appendix B.
5 One-Dimensional CSBM
In Baranwal et al. (2021), the authors showed that analyzing the linear classifier for the $m$-dimensional CSBM reduces to analyzing the $1$-dimensional model. We say that a CSBM is one-dimensional and centered with parameters $n,p,q,\sigma$ if it has one-dimensional features and means $1/\sqrt{n}$ and $-1/\sqrt{n}$. That is, we have one feature vector $x\in\mathbb{R}^{n}$ given by $x=s+g$, where $g\sim\mathcal{N}(0,\sigma^{2}I_{n})$ and $s(i)=1/\sqrt{n}$ for $i\in S$ and $-1/\sqrt{n}$ for $i\in T$. We will refer to $s$ as our signal vector and for ease of notation, we normalize it so that it always has unit norm. The following lemma allows us to reduce the analysis of the linear classifier for a general CSBM to the analysis of the 1-dimensional centered CSBM (proof in Section B.1). Thus, in the proofs of our main theorems, we will analyze the 1-dimensional case before applying Lemma 5.1.
Lemma 5.1.
Given an $m$-dimensional 2-block $CSBM$, there exist $w\in\mathbb{R}^{m}$ and $b\in\mathbb{R}^{n}$ such that $Xw+b=s+g$ where for each vertex $i$, $g_{i}$ is i.i.d. $\mathcal{N}(0,\sigma^{\prime 2})$ with $\sigma^{\prime}=\frac{4\sigma}{\sqrt{n}\left\lVert\mu-\nu\right\rVert}$.
In the 1-dimensional model, it is clear how our signal $s$ is present in our features. Our convolution matrix also captures the signal because it can be viewed as a perturbation of the matrix $ss^{\top}$. This is especially evident with the un-normalized convolution matrix $\tilde{A}$, which satisfies the following
$\tilde{A}=\eta ss^{\top}+\frac{1}{d}R+d^{\prime}\mathbbm{1}\mathbbm{1}^{\top}$ (2)
where $\eta:={(p-q)n}/({2d})$ is the signal strength, $d^{\prime}:=({(p+q)n/2-d})/({nd})$ is the average degree deviation, and $R$ is the “edge-deviation" matrix, where $R_{i,j}=A_{i,j}-\mathbb{E}[A_{i,j}]$. Since $R$ has i.i.d. zero-mean entries with variance at most $p$, we can use standard matrix concentration inequalities to show it is not too big. Likewise, $d^{\prime}$ is small due to degree concentration, and together, these two concentration results imply that $\tilde{A}$ is close to $\gamma ss^{\top}$. In fact, if we show degree concentration for all vertices, then we can show that $\hat{A}$ also behaves like $\tilde{A}$. We state these concentration results below.
Proposition 5.2.
Assume that $p=\Omega(\frac{\log^{2}{n}}{n})$, and let $\gamma=\frac{p-q}{p+q}$. With probability $1-n^{-\Omega(1)}$, we have the following concentration results
  1. 1.
    $|d^{\prime}|\leq O(1/n^{1.5})$, which implies that $\eta\in\gamma(1\pm o(1))$.
  2. 2.
    $\left\lVert R\right\rVert\leq O(\sqrt{np})$
  3. 3.
    $\left\lVert\tilde{A}-\gamma ss^{\top}\right\rVert\leq O(\frac{1}{\sqrt{np}})$ and $\left\lVert\hat{A}-\gamma ss^{\top}\right\rVert\leq O(\sqrt{\frac{\log{n}}{np}})$
These concentration properties are crucial for spectral analysis. Details are given in Appendix B.
6 Details about our Partial Classification Result
In this section, we give a sketch of the proof of our partial classification result, Theorem 4.1. We will show that partial classification can be achieved if the convolved vector is well correlated with our signal vector $s$ as defined in the beginning of Section 5 in the centered-1 dimensional CSBM. We will analyze our result for convolutions using the matrix $M\in\{\tilde{A},\hat{A}\}$.
Proposition 6.1.
Given a centered 1-dimensional 2 block CSBM with parameters $n,p>q,\;\sigma$ and $\gamma(p,q)=\frac{p-q}{p+q}$, suppose our convolution matrix $M$ satisfies $\left\lVert M-\gamma ss^{\top}\right\rVert\leq\delta$ and $\gamma\geq C\delta$ for a large enough constant $C$. Let $x_{k}=M^{k}x$, be the result of applying $k$ rounds of graph convolution to our input feature $x$. Then with probability at least $1-\frac{1}{2}\exp(-\frac{1}{4\sigma^{2}})$, there exists a scalar $C_{k}$ and an absolute constant $C^{\prime}$ such that
$\left\lVert C_{k}x_{k}-s\right\rVert^{2}\leq O\left(\frac{\delta^{2}}{\gamma^{2}}+\Big{(}\frac{C^{\prime}\delta}{\gamma}\Big{)}^{2k}n\sigma^{2}\log{n}\right)$
The main idea of our analysis is to use the fact that the top eigenvector of $M$, denoted $\hat{s}$, is well correlated with our signal $s$. Since $\left\lVert M-\gamma ss^{\top}\right\rVert\leq\delta$ and $\gamma>C\delta$ by assumption, standard Matrix perturbation arguments imply that the spectrum of $M$ will be in $(\gamma\pm\delta,\pm\delta,\pm\delta,...)$ with high probability. Given our assumption of $\gamma>C\delta$, there will be a large gap between the top eigenvalue of $M$ and the rest of its eigenvalues. A well-known result Davis and Kahan implies that $\left\lVert s-\hat{s}\right\rVert^{2}\leq O(\delta^{2}/\gamma^{2})$, i.e. $s$ is close to $\hat{s}$. Thus, we prove Proposition 6.1 by showing that the influence of the rest of the eigenvectors on our convolved vector, $x_{k}=M^{k}x$, decreases exponentially with $k$, which allows us to bound the squared norm distance between $x_{k}$ and $s$. In particular, we take our normalization constant to be $C_{k}\approx 1/\lambda_{1}^{k}$, where $\lambda_{1}$ is the maximum eigenvalue of $M$. Note that $\lim_{k\rightarrow\infty}({1}/{\lambda_{1}^{k}})M^{k}=\hat{s}\hat{s}^{\top}$. Roughly speaking, we decompose our convolution vector as $C_{k}x_{k}\approx({1}/{\lambda_{1}^{k}})M^{k}s+({1}/{\lambda_{1}^{k}})M^{k}g$. To bound the distance of this vector from $s$, we analyze the contribution of each of the two terms to our error separately. That is, we show that with high probability $\|({1}/{\lambda_{1}^{k}})M^{k}s-s\|^{2}\leq O(\delta^{2}/\gamma^{2})$ and $\left\lVert M^{k}g\right\rVert^{2}\leq O((\delta/\gamma)^{2k}n\sigma^{2}\log{n})$. Note that the first error term is from taking the convolution of the noisy graph with the true signal, and thus does not decrease with $k$. The second error term, on the other hand, comes from variance in the features, $g$, and thus decreases with our noise level $\sigma$ and drops exponentially with each convolution.
Finally, given Proposition 6.1, we can prove the partial classification result by noting that if we partition the convolved 1-dimensional data around 00, then each misclassified vertex contributes $1/n$ to the mean-squared error, which means the number of misclassified vertices is at most $\left\lVert C_{k}x_{k}-s\right\rVert^{2}n$. This, combined with Lemma 5.1 to generalize to the $m-$dimensional case will prove Theorem 4.1
7 Details about our Exact Classification Result
In this section, we sketch the proof of Theorem 4.2 for exact classification using the un-normalized corrected convolution matrix $\tilde{A}$. To show linear separability, we would like $x_{k}=\tilde{A}^{k}x$ to have positive entries for all vertices in $S$ and negative entries for all vertices in $T$. This means that we want to show $\left\lVert C_{k}x_{k}-s\right\rVert_{\infty}<1/\sqrt{n}$ for some appropriate scalar $C_{k}$. In particular, we will take $C_{k}$ to be $1/\eta^{k}$, where $\eta$ is our empirical estimate of $\gamma(p,q)$. In partial classification, it sufficed to bound the mean squared error $\left\lVert C_{k}x_{k}-s\right\rVert_{2}^{2}$, using spectral analysis but bounding $\left\lVert C_{k}x_{k}-s\right\rVert_{\infty}$ requires more work because now we are bounding the entrywise instead of average error. In our approach, we bound the volume of messages passed through “incorrect paths” in our graph and show that the contribution from these messages is small. Then, we show the other source of error, the feature variance, is reduced exponentially with each round of convolution. As with the partial classification result, we first prove our result in the 1-dimensional centered model:
Proposition 7.1.
Suppose we are given a 1-dimensional centered 2-block CSBM with parameters $n,p,q,\sigma$ and $k=O(\log{n})$ such that $\gamma=\frac{p-q}{p+q}\geq\Omega\left(k\sqrt{\frac{\log{n}}{np}}\right)$, $p\geq\frac{\log^{3}{n}}{n}$, and $\sigma\leq O\left(\frac{1}{\sqrt{\log{n}}}\right)$. Then with probability at least $1-n^{-\Omega(1)}$:
$\left\lVert\frac{1}{\eta^{k}}\tilde{A}^{k}x-s\right\rVert_{\infty}\leq\frac{1}{2\sqrt{n}}+O\left(\left(\frac{C}{\gamma\sqrt{np}}\right)^{k}\sigma\sqrt{\log{n}}\right)$
Given, Proposition 7.1, Theorem 4.2 follows immediately by applying Lemma 5.1 as with the proof of Theorem 4.1. We now give a sketch of our proof for Proposition 7.1. To bound each entry of $C_{k}x_{k}-s$, we must bound $|e_{u}^{\top}\tilde{A}^{k}x/\eta^{k}-e_{u}^{\top}s|$ for all $u\in V$. Similar to the partial classification case, we will split our error into error from $\tilde{A}^{k}s$ and error from $\tilde{A}^{k}g$. That is for each $u\in V$, we upper bound $|e_{u}^{\top}(\eta ss^{\top}+R^{\prime})^{k}s-e_{u}^{\top}s|$ and $|e_{u}^{\top}(\eta ss^{\top}+R^{\prime})^{k}g|$, where $R^{\prime}=\tilde{A}-\eta ss^{\top}$, and we show both are small with high probability. The matrix $(\eta ss^{\top}+R^{\prime})^{k}$, when expanded out, can be written as $\eta^{k}ss^{\top}$ plus a sum of $12^{k}-$ 1 terms, each of which is a non-commutative product of matrices of the form $\eta ss^{\top}$ or $R^{\prime}$. We group these error matrices into terms of order $\ell$ for $\ell\in[k]$, where the $\ell^{th}$ order terms are comprised of products that contain $\ell$ copies of $R^{\prime}$ and $k-\ell$ copies of $\gamma ss^{\top}$.
Our analysis consists of three steps. First, we use degree concentration to show that instead of analyzing the error w.r.t. $R^{\prime}$, it suffices to analyze the error w.r.t. $\frac{1}{d}R$. Second, we control the term $e_{u}^{\top}R^{\ell}s$ for any $\ell\in[k]$. In particular, we show that with high probability, for all $u\in V$ and $\ell\in[k]$, $|e_{u}^{\top}R^{\ell}s|\leq{(Cnp\log{n})^{\ell/2}}/{\sqrt{n}}$ for a constant $C$. This part requires the slightly stronger graph density assumption: $p\geq\Omega({\log^{3}{n}}/{n})$. Finally, to prove the main proposition, we show that after expanding out $e_{u}^{\top}\tilde{A}^{k}s$, each error term arising from an error matrix of order $\ell$ can be written as $e_{u}^{\top}R^{a_{1}}s\cdot s^{\top}R^{a_{2}}s\cdot...s^{\top}R^{a_{L}}s$ where $a_{1},...a_{L}$ are non-negative integers satisfying $a_{1}+a_{2}+...a_{L}=\ell$. We can apply what we proved in part 2 to bound the first term and for the rest of the terms, we use $|s^{\top}R^{a}s|\leq\left\lVert R\right\rVert^{a}$ and with high probability $\left\lVert R\right\rVert\leq O(\sqrt{np})$. Thus, we have $|e_{u}^{\top}R^{a_{1}}s\cdot s^{\top}R^{a_{2}}s\cdot...s^{\top}R^{a_{L}}s|\leq\frac{1}{\sqrt{n}}{(Cn\log{n})^{\ell/2}}$ for some constant $C$. The analysis for bounding $e_{u}\tilde{A}^{k}g$ is similar. Finally, by combining these bounds with our assumptions that $\gamma$ is large enough, we obtain Proposition 7.1.
Now, we take a closer look at the second step of bounding $e_{u}^{\top}R^{\ell}s$, which is the main technical part of this analysis. Recall that $R$ is a random symmetric matrix with i.i.d. zero mean entries. The term $e_{u}^{\top}R^{\ell}s$ can be expressed as $\sum_{w}R_{w(0),w(1)}R_{w(1),w(2)},...R_{w(\ell-1),w(\ell)}s(w(\ell))$ where the sum is over all walks, $w$, of length $\ell$ in the complete graph over $n$ vertices starting at $w(0):=u$. By upper bounding this term, we are upper-bounding the total amount of message $u$ can receive through “incorrect paths” of length $\ell$ in the graph. In the case of $\ell=1$, we can think of this as bounding the difference between the amount of message $u$ receives from its neighbors and the amount of message it expects to receive, which will always have more from the intra-class neighbors. To show that $|e_{u}^{\top}R^{\ell}s|$ is bounded with high probability, we use a path counting argument to control the $2t^{th}$ moment of this random variable, for $t=O(\log{n}/\ell)$. Then, we apply the standard Markov’s inequality $\mathbf{Pr}\left[|e_{u}^{\top}R^{\ell}s|\geq\lambda\right]<\mathbb{E}[|e_{u}^{\top}R^{\ell}s|^{2t}]/\lambda^{2t}$. Our argument is inspired by the proof of the central limit theorem using pairing partitions, as in chapter 8 of Nica and Speicher (2010).
8 Experiments
In this section, we demonstrate our results empirically. For synthetic data, we show Theorems 4.1 and 4.2 for linear binary classification. For real data, we show that removing the principal component of the adjacency matrix exhibits positive effects on multi-class node classification problems as well.
8.1 Synthetic Data
For synthetic data from the CSBM, we demonstrate the benefits of removing the principal component of the adjacency matrix before performing convolutions for both variants of convolution described in Equation 1. We choose $n=2000$ nodes with $20$ features for each node, sampled from a Gaussian mixture. The intra-edge probability is fixed to $p=O(\log^{3}n/n)$. We perform linear classification to demonstrate the results in Theorem 4.1 and Theorem 4.2, training a one-layer GCN network both with and without the corrected convolutions and perform an empirical comparison.
We provide plots for two different settings: (1) Fix $\gamma={|p-q|}/({p+q})=2/3$ and vary signal-to-noise ratio of the node features, $\|\mu-\nu\|/\sigma$, for different number of convolutions. We observe in Figure 1 that as the number of convolutions increases, the original GCN (Kipf and Welling, 2017) (in blue) starts performing poorly, while the corrected versions (in orange and green) retain the accuracy for lower signal-to-noise ratio; (2) Fix ${\|\mu-\nu\|}/{\sigma}=1$ and vary the graph relative signal strength, $\gamma$, for different number of convolutions. We observe the same trends in this setting, as depicted in Figure 2. The vertical lines represent the threshold for exact classification from Theorem 4.2.
(a) $1$ convolution.
(b) $2$ convolutions.
(c) $4$ convolutions.
(d) $8$ convolutions.
(e) $12$ convolutions.
(f) $16$ convolutions.
Figure 1: Accuracy plot (average over 50 trials) against the signal-to-noise ratio of the features (ratio of the distance between the means to the standard deviation) for increasing number of convolutions. Here, $v=D^{1/2}\mathbbm{1}$ and the “GCN with $vv^{\top}$ removed” refers to convolution with the corrected, normalized adjacency matrix. “GCN with $\mathbbm{1}\mathbbm{1}^{\top}$ removed” is the corrected, unnormalized matrix.
(a) $1$ convolution.
(b) $2$ convolutions.
(c) $4$ convolutions.
(d) $8$ convolutions.
(e) $12$ convolutions.
(f) $16$ convolutions.
Figure 2: Accuracy plot (average over 50 trials) against graph relative signal strength ($\gamma=|p-q|/(p+q)$) for various values of the number of convolutions.
8.2 Real Data
Similar to synthetic data, we compare the results for corrected graph convolution to the original GCN on the following real graph benchmarks datasets: CORA, CiteSeer, and Pubmed citation networks (Sen et al., 2008). It is observed in Figure 3 that correcting the convolution as in Equation 1 by removing the principal component performs better as the number of layers increases, even for a multi-class setting, where we perform one convolution at each layer.
Note that in Figure 3, we see that overall the accuracy of every learning method decreases as the number of convolutions increases but the corrected convolutions converge to an accuracy much higher than that of the uncorrected convolution. This is attributed to the fact that for multi-class classification, the important information about class memberships is captured by the top $C$ eigenvectors (except the first one) where $C$ is the number of classes (Lee et al., 2014). Since the limiting behavior of many rounds of convolutions is akin to projecting the features onto the second eigenvector, we only expect this to capture partial information about the multi-class structure.
(a) Cora.
(b) CiteSeer.
(c) PubMed.
Figure 3: Accuracy plots (average over 50 trials) against the number of layers for real datasets.
9 Conclusion and Future Work
In this study, we utilized spectral methods to obtain partial and exact classification results for the linear classifier with corrected convolution matrices in the 2-block CSBM. Our spectral approach highlights, theoretically, how removing the top eigenvector can mitigate oversmoothing and improve classification accuracy. We prove that the removal of the top eigenvector results in reducing feature variance and correcting the asymptotic behavior of many rounds of convolution towards the second, rather than the top eigenvector of the adjacency matrix. We hope our analysis can lead to further developments in theoretical and practical studies of GNNs. On the theoretical side, we believe our analysis can be extended to more general models. One potential avenue would be to analyze the multi-class CSBM using more sophisticated architectures, like the spectral convolutional network with low-degree matrix polynomials (Defferrard et al., 2016) and non-linear activation functions.
Acknowledgements
K. Fountoulakis would like to acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC). Cette recherche a été financée par le Conseil de recherches en sciences naturelles et en génie du Canada (CRSNG), [RGPIN-2019-04067, DGECR-2019-00147].
Appendix A Linear Algebra and Probability Background
Before beginning our proofs, we will establish the following theorems from the literature and basic facts that we will use throughout the proofs.
A.1 Linear Algebra
For matrix inequalities, we use the following inequalities about matrix spectral norms.
Theorem A.1.
([Vershynin, 2018] theorem 4.5.3.) Let $A$ be a symmetric matrix with eigenvalues $\lambda_{1}\geq\lambda_{2}\geq...\lambda_{n}$ and $B$ be a symmetric matrix $\mu_{1}\geq\mu_{2}\geq....\mu_{n}$. Suppose $\left\lVert A-B\right\rVert\leq\delta$. Then, $\max_{i}\left\lVert\lambda_{i}-\mu_{i}\right\rVert\leq\delta$
We also use the following basic result to relate the matrix spectral norm to its maximum entry
Lemma A.2.
Let $M$ be an $n\times n$ symmetric matrix such that each row of $M$ has at most $m$ non-zero entries and each entry of $M$ has an absolute value at most $\varepsilon$. Then $\left\lVert M\right\rVert\leq\varepsilon m$
Proof.
We will use the fact that for any scalars $a,b$, we have $2ab\leq a^{2}+b^{2}$ since $a^{2}+b^{2}-2ab=(a-b)^{2}\geq 0$. We will also use $\text{supp}(M)$ to denote the set of non-zero entries in $M$. Let $x$ be a unit vector. Then we have
$|x^{\top}Mx|$ $=|\sum_{i}\sum_{j=1}^{n}x(i)x(j)M_{i,j}|$
$\leq\varepsilon\sum_{i,j\in\text{supp}(M)}x(i)x(j)$
$\leq\varepsilon\sum_{i,j\in\text{supp}(M)}\frac{1}{2}x(i)^{2}+\frac{1}{2}x(j)^{2}$
$\leq\varepsilon m\sum_{i}x(i)^{2}$
$=\varepsilon m$
Where the last inequality follows from the fact for every $i$, there are at most $m$ entries $j$ such that $M_{i,j}\neq 0$. ∎
A.2 Concentration Inequalities
Throughout our analysis, we will mainly use the following two concentration inequalities. The first is a bound on the the expected deviation of the sum of Bernoulli random variables
Theorem A.3.
Let $X_{1},...X_{n}$ be Bernoulli random variables with mean at most $p$ and $S_{n}=\sum_{i=1}^{n}X_{i}$. Then
$\mathbf{Pr}\left[|S_{n}-\mathbb{E}[S_{n}]|>t\right]<\exp(-\Omega(\frac{t^{2}}{np}))$
The second concentration inequality we will upper-bounds the spectral norm of random matrices whose entries have zero mean and bounded variance.
Theorem A.4.
([Bandeira and van Handel, 2016] Remark 3.13) Let $R$ be a random matrix whose entries $R_{i,j}$ are independent, zero mean, with variance $\sigma^{2}$. Then with probability $1-n^{-\Omega(1)}$, we have
$\left\lVert R\right\rVert\leq O(\sigma\sqrt{n}+\log{n})$
In particular, if $\sigma^{2}=\Omega(\frac{\log^{2}{n}}{n})$ then $\left\lVert R\right\rVert\leq O(\sigma\sqrt{n})$
We will also use the standard tail bound on the norm of a Gaussian vector
Lemma A.5.
Let $g\sim\mathcal{N}(0,\sigma^{2}I)$ be a random Gaussian vector. Then $\mathbf{Pr}\left[\left\lVert g\right\rVert>t\right]\leq 2\exp(-\frac{t^{2}}{2n\sigma^{2}})$
A.3 Other Facts
We will also be using the following basic fact about approximating the exponential function
Lemma A.6.
For all $x$, $1+x\leq e^{x}$. If $x\leq 1$, then $e^{x}\leq 1+2x$.
Appendix B Proofs in Section 5
B.1 Proof of Lemma 5.1
In the following section, we give the missing proofs of the basic properties of the CSBM. The first is the reduction from the $m-$dimensional model to the centered $1-$dimensional model, which we will restate here:
Lemma B.1.
Given an $m$-dimensional 2-block $CSBM$, there exist $w\in\mathbb{R}^{m}$ and $b\in\mathbb{R}^{n}$ such that $Xw+b=s+g$ where for each vertex $i$, $g_{i}$ is i.i.d. $\mathcal{N}(0,\sigma^{\prime 2})$ with $\sigma^{\prime}=\frac{4\sigma}{\sqrt{n}\left\lVert\mu-\nu\right\rVert}$.
Proof.
Let $X_{i,:}$ be the $i^{th}$ row of $X$. For each $i\in S$, we have $X_{i,:}=\mu+g_{i}$ and for each $j\in T$ we have $X_{j,:}=\nu+g_{j}$, where $g_{i}\sim\mathcal{N}(0,\sigma^{2}I_{m})$ are i.i.d. random Gaussian noise vectors. Now we pick $w=(\mu-\nu)$, $b=-\frac{1}{2}\langle\mu+\nu,\mu-\nu\rangle\vec{\mathbbm{1}}$. Now $x^{\prime}=Xw+b$. For $i\in S$, we let
$s^{\prime}(i):=\langle\mu-\frac{1}{2}(\mu+\nu),\mu-\nu\rangle=\frac{1}{2}\left\lVert\mu-\nu\right\rVert^{2}$
and for $i\in T$, we let
$s^{\prime}(i):=\langle\nu-\frac{1}{2}(\mu+\nu),\mu-\nu\rangle=-\frac{1}{2}\left\lVert\mu-\nu\right\rVert^{2}$
The $s^{\prime}$ part is our signal. On the other hand, the noise at each entry is given by $x^{\prime}(i)-s^{\prime}(i)=\langle g_{i},\mu-\nu\rangle$ which is Gaussian with standard deviation $\sigma\left\lVert\mu-\nu\right\rVert$. Now, if we let $x=\frac{4}{\sqrt{n}\left\lVert\mu-\nu\right\rVert^{2}}x^{\prime}$, then we have $x=s+g$, where $s(i)\in\pm\frac{1}{\sqrt{n}}$ and each entry of $g$ is i.i.d. Gaussian with standard deviation $\sigma^{\prime}:=\frac{4\sigma}{\sqrt{n}\left\lVert\nu-\mu\right\rVert}$. ∎
Proof of Proposition 5.2 and Characterizing the Convolution Matrix
In this section, we will provide some crucial properties of the graph convolution matrices in (1) and then use them to prove Proposition 5.2. The adjacency matrix $A$ is a random matrix with expectation
$\mathbb{E}[A]=\begin{bmatrix}pJ_{n/2}&qJ_{n/2}\\ qJ_{n/2}&pJ_{n/2}\end{bmatrix}$
Where $J_{n/2}$ is $n/2\times n/2$ all-ones matrix. It can be seen that $\mathbb{E}[A]$ is a rank-2 matrix, with eigenvectors $\frac{1}{\sqrt{n}}\mathbbm{1}$ and $s$ because $\mathbb{E}[A]\mathbbm{1}=\frac{1}{2}(p+q)n\mathbbm{1}$ and $\mathbb{E}[A]s=\frac{1}{2}(p-q)ns$, which means
$\mathbb{E}[A]=\frac{1}{2}(p+q)\mathbbm{1}\mathbbm{1}^{\top}+\frac{1}{2}(p-q)nss^{\top}$
Thus, in the centered one-dimensional CSBM, our true signal, $s$, is encoded in two ways: as the second eigenvector of the matrix $\mathbb{E}[A]$ and as $\mathbb{E}[x]$. We can express our unnormalized corrected convolution matrix in terms of the signal, $ss^{\top}$, as follows:
$\tilde{A}$ $=\frac{1}{d}A-\frac{1}{n}\mathbbm{1}\mathbbm{1}^{\top}=\frac{(p-q)n}{2d}ss^{\top}+\frac{\frac{1}{2}(p+q)n-d}{nd}\mathbbm{1}\mathbbm{1}^{\top}+\frac{1}{d}(A-\mathbb{E}[A])$
Recall from main text that we defined $\eta:=\frac{(p-q)n}{2d}$ be the signal strength, $d^{\prime}:=\frac{(p+q)n/2-d}{nd}$ to be the degree deviation, and $R=A-\mathbb{E}[A]$ to be the error matrix. Thus, we have
$\tilde{A}=\eta ss^{\top}+\frac{1}{d}R+d^{\prime}\mathbbm{1}\mathbbm{1}^{\top}$
Note that $R$ has independent zero mean entries with bounded variance, and such random matrices have well-studied concentration properties. The main idea of our analysis is to show that $d^{\prime}$ and $R$ are small compared to our matrix signal strength $\eta$ with high probability so that $\tilde{A}$ behaves like $\eta ss^{\top}\approx\gamma ss^{\top}$. To analyze convolution with the normalized $\hat{A}$, we use degree-concentration to show that $\hat{A}$ is close to $\tilde{A}$, and so also behaves like $\gamma ss^{\top}$.
Now we prove the main proposition in our section, Proposition 5.2, which we restate as follows:
Proposition B.2.
Assume that $p=\Omega(\frac{\log{n}}{n})$, and let $\gamma=\frac{p-q}{p+q}$. With probability $1-n^{-\Omega(1)}$, we have the following concentration results
  1. 1.
    $|d^{\prime}|=\frac{|d-\frac{1}{2}(p+q)n|}{nd}\leq O(1/n^{1.5})$, which implies that $\eta\in\frac{p-q}{p+q}(1\pm o(1))$.
  2. 2.
    $\left\lVert R\right\rVert\leq O(\sqrt{np})$
  3. 3.
    $\left\lVert\tilde{A}-\gamma ss^{\top}\right\rVert\leq O(\frac{1}{\sqrt{np}})$ and $\left\lVert\hat{A}-\gamma ss^{\top}\right\rVert\leq O(\sqrt{\frac{\log{n}}{np}})$
Proof.
To bound the deviation of the average degree is equivalent to bounding the total number of edges in the graph. Note that the expected number of edges in the graph is $\frac{1}{4}(p+q)n^{2}$. Thus, we apply Theorem A.3 with $t=\Theta((p+q)n^{1.5})$ to obtain
$\mathbf{Pr}\left[||E|-\frac{1}{4}(p+q)n^{2}|>t\right]\leq\exp(-\Omega(pn))=n^{-\Omega(1)}$
Since $d=2|E|/n$, we have that with high probability, $|d-\frac{1}{2}(p+q)n|\leq O((p+q)\sqrt{n})$ and $|d^{\prime}|=\frac{|d-\frac{1}{2}(p+q)n|}{nd}\leq O(\frac{(p+q)\sqrt{n}}{(p+q)n^{2}(1-o(1)))}\leq O(1/n^{1.5})$. This gives us bound number 1. To get bound number 2, we apply Theorem A.4 on the error matrix $R$, noting that each entry of $R$ has variance at most $p$. To get bound number 3, let $R^{\prime}=\frac{1}{d}R+d^{\prime}\mathbbm{1}\mathbbm{1}^{\top}=\tilde{A}-\eta ss^{\top}$. Then with high probability, we have
$\left\lVert R^{\prime}\right\rVert$ $\leq\frac{1}{d}\left\lVert R\right\rVert+d^{\prime}\left\lVert J\right\rVert$
$\leq\frac{1}{\frac{1}{2}(p+q)n(1-o(1))}\left\lVert R\right\rVert+d^{\prime}\left\lVert\mathbbm{1}\mathbbm{1}^{\top}\right\rVert$
$\leq O(\frac{1}{\sqrt{pn}}+\frac{1}{\sqrt{n}})\quad\text{since $\left\lVert\mathbbm{1}\mathbbm{1}^{\top}\right\rVert=n$}$
$=O(\frac{1}{\sqrt{pn}})$
Finally, we analyze the corrected normalized adjacency matrix $\hat{A}$. In this case, we want to show degree concentration for every node. Let $d_{v}$ be the degree of node $v$, and $\bar{d}=\frac{1}{2}(p+q)n$ be the expected degree. By applying Theorem A.3 and using our assumption $p>q$, we have
$\mathbf{Pr}\left[|d_{v}-\bar{d}|>\sqrt{Cnp\log{n}}\right]\leq\exp(-\Omega(\log{n}))$
Thus, with high probability, we have $|d_{v}-d|\leq O(\sqrt{np\log{n}})$ for all $v\in V$. Now, to analyze the normalized adjacency matrix, we have
$\left\lVert\hat{A}-\gamma ss^{\top}\right\rVert$ $=\left\lVert\hat{A}-\gamma ss^{\top}+\tilde{A}-\tilde{A}\right\rVert$
$\leq\left\lVert\hat{A}-\tilde{A}\right\rVert+\left\lVert\tilde{A}-\gamma ss^{\top}\right\rVert$
$\leq\left\lVert D^{-1/2}AD^{-1/2}-\frac{1}{d}A\right\rVert+\left\lVert\frac{1}{\sum_{v}d_{v}}D^{1/2}\mathbbm{1}\mathbbm{1}^{\top}D^{1/2}-\frac{1}{n}\mathbbm{1}\mathbbm{1}^{\top}\right\rVert+O(\frac{1}{\sqrt{np}})$
To bound $\left\lVert D^{-1/2}AD^{-1/2}-\frac{1}{d}A\right\rVert$, let $\varepsilon=\sqrt{\frac{C\log{n}}{np}}$ for large enough $C$ so that for all $v$, $d_{v}\in\bar{d}(1\pm\varepsilon)$. We note that the adjacency matrix has at most $\bar{d}(1+\varepsilon)$ non-zero entries for each row and for each $u,v$, we have
$|\frac{1}{\sqrt{d_{u}d_{v}}}-\frac{1}{d}|\leq\frac{1}{\bar{d}(1-\varepsilon)}-\frac{1}{\bar{d}(1+\varepsilon)}\leq O(\frac{\varepsilon}{d})$
Thus, by applying Lemma A.2, we have $\left\lVert D^{-1/2}AD^{-1/2}-\frac{1}{d}A\right\rVert\leq O(\varepsilon)$. Similarly, for all $u,v$, we have
$|\frac{\sqrt{d_{u}d_{v}}}{\sum_{i}d_{i}}-\frac{1}{n}|\leq\frac{\bar{d}(1+\varepsilon)}{\bar{d}n(1-\varepsilon)}-\frac{1}{n}\leq O(\frac{\varepsilon}{n})$
Thus, Lemma A.2 implies that $\left\lVert\frac{1}{\sum_{v}d_{v}}D^{1/2}\mathbbm{1}\mathbbm{1}^{\top}D^{1/2}-\frac{1}{n}\mathbbm{1}\mathbbm{1}^{\top}\right\rVert\leq O(\varepsilon)$. Finally, this implies that
$\left\lVert\hat{A}-\gamma ss^{\top}\right\rVert\leq O(\sqrt{\frac{\log{n}}{np}})$
Note that the case $p>q$ corresponds to a homophilous graph, and the case $p<q$ corresponds to a heterophilous graph (see Luan et al. [2021], Ma et al. [2022] for more), however, for binary classification, it has been shown [Baranwal et al., 2023a] that once can assume $p>q$ without loss of generality, and make corresponding adjustments in the classifier. Indeed in our case, if $p<q$ then we can take $-\tilde{A}$ as our convolution matrix so that its maximum modulus eigenvalue is positive. Moreover, since our signal, $s$, is centered, taking a convolution with $\tilde{A}$ is equivalent to taking a convolution with $-\tilde{A}$. Thus, we can assume without loss of generality that $\gamma$ is always non-negative, i.e. $p>q$ for the rest of our analysis.
Appendix C Proofs in Section 6
In this section, we will formally prove our first main theorem on partial classification, Theorem 4.1. To start, we will show the correlation between the top eigenvector of the convolution matrix $M$ and $s$. Note that this result is well-known in the literature (for example see Vershynin [2018] chapter 4.5). We will give the proof here for completeness.
Lemma C.1.
Suppose $M$ satisfies $\left\lVert M-\gamma ss^{\top}\right\rVert\leq\delta$ where $\gamma>C\delta$ for large enough constant $C$. Let $\hat{s}$ be the top eigenvector of $M$. Then $\langle s,\hat{s}\rangle^{2}\geq 1-4\frac{\delta^{2}}{\gamma^{2}}$, and $\left\lVert s-\hat{s}\right\rVert^{2}=O(\delta/\gamma)$.
Proof.
By assumption, we can express $M$ as $\gamma ss^{\top}+R^{\prime}$ where $\left\lVert R^{\prime}\right\rVert\leq\delta$. Now let $\lambda_{k}$ be the $k^{th}$ largest eigenvalue of $M$. By Theorem A.1, we have $|\lambda_{1}-\gamma|\leq\delta$ and $\forall i>1$, $|\lambda_{i}|\leq\delta$. Now consider the matrix $I-ss^{\top}$, which is the projection matrix onto the orthogonal complement subspace of $s$. Since $M=\gamma ss^{\top}+R^{\prime}$, we have
$\gamma(I-ss^{\top})=\gamma I-M+R^{\prime}$
and the matrix $\gamma I-M$ has the same eigenvectors as $M$. In particular, $\hat{s}$ is an eigenvector with eigenvalues $\gamma-\lambda_{1}$. Thus, we have
$\gamma\sqrt{1-\langle s,\hat{s}\rangle^{2}}$ $=\gamma\sqrt{\hat{s}^{\top}(I-ss^{\top})\hat{s}}$
$=\left\lVert\gamma(I-ss^{\top})\hat{s}\right\rVert$
$=\left\lVert(\gamma I-M)\hat{s}+R^{\prime}\hat{s}\right\rVert$
$\leq\left\lVert(\gamma I-M)\hat{s}\right\rVert+\left\lVert R^{\prime}\hat{s}\right\rVert$
$\leq|\gamma-\lambda_{1}|+\delta$
$\leq 2\delta$
Thus, we have $\langle s,\hat{s}\rangle^{2}\geq 1-4\delta^{2}/\gamma^{2}$. Since by our assumption, $\delta/\gamma$ is small, $\hat{s}$ is very close to either $s$ or $-s$. For our analysis, it doesn’t matter which of these is the case as both $s$ and $-s$ equally separate the points in our two classes. Thus, we can assume WLOG that $\langle s,\hat{s}\rangle>0$. As a consequence, we have
$\left\lVert s-\hat{s}\right\rVert^{2}$ $=2-2\langle s,\hat{s}\rangle=2-2\sqrt{1-2\delta/\gamma}=O(\delta/\gamma)$
where the second inequality follows from applying the first order approximation $\sqrt{1-x}=1-\frac{x}{2}-O(x^{2})$ using the fact that $\delta/\gamma$ is small. Note that if $\langle s,\hat{s}\rangle$ was negative, we can do the same analysis in the following sections with $-s$ instead of $s$. ∎
Now we will prove the main result about partial classification in the $1$-dimensional model, Proposition 6.1, which we restate as follows:
Proposition C.2.
Given a centered 1-dimensional 2 block CSBM with parameters $n,p>q,\;\sigma$ and $\gamma(p,q)=\frac{p-q}{p+q}$, suppose our convolution matrix $M$ satisfies $\left\lVert M-\gamma ss^{\top}\right\rVert\leq\delta$ and $\gamma\geq C\delta$ for a large enough constant $C$. Let $x_{k}=M^{k}x$, be the result of applying $k$ rounds of graph convolution to our input feature $x$. Then with probability at least $1-\frac{1}{2}\exp(-\frac{1}{4\sigma^{2}})$, there exists a scalar $C_{k}$ and an absolute constant $C^{\prime}$ such that
$\left\lVert C_{k}x_{k}-s\right\rVert^{2}\leq O\left(\frac{\delta^{2}}{\gamma^{2}}+\Big{(}\frac{C^{\prime}\delta}{\gamma}\Big{)}^{2k}n\sigma^{2}\log{n}\right)$
Proof.
Let $\lambda_{1}\geq...\lambda_{n}$ be the eigenvalues of $M$ and $w_{1}=:\hat{s},w_{2}...w_{n}$ be their corresponding eigenvectors. By Lemma C.1, the eigenvalues satisfy $\lambda_{1}\in\gamma\pm\delta$ and $|\lambda_{i}|\leq\delta$ for $i>1$. The convolution vector after $k$ rounds can be expressed as $x_{k}=M^{k}g+M^{k}s$, and we will analyze each term individually.
$M^{k}g$ $=\lambda_{1}^{k}\langle g,w_{1}\rangle\hat{s}+\sum_{i>1}\lambda_{i}^{k}\langle g,w_{i}\rangle w_{i}$
$=\lambda_{1}^{k}\langle g,w_{1}\rangle s+\lambda_{1}^{k}\langle g,w_{1}\rangle(\hat{s}-s)+\sum_{i>1}\lambda_{i}^{k}\langle g,w_{i}\rangle w_{i}$
and similarly,
$M^{k}s$ $=\lambda_{1}^{k}\langle s,w_{1}\rangle\hat{s}+\sum_{i>1}\lambda_{i}^{k}\langle s,w_{i}\rangle w_{i}$
$=\lambda_{1}^{k}\langle s,w_{1}\rangle s+\lambda_{1}^{k}\langle s,w_{1}\rangle(\hat{s}-s)+\sum_{i>1}\lambda_{i}^{k}\langle s,w_{i}\rangle w_{i}$
Now we will let $E_{g}$ and $E_{s}$ be the error vectors, i.e. vectors that are not in the span of $s$, from each of the terms respectively. That is
$E_{g}$ $=\lambda_{1}^{k}\langle g,w_{1}\rangle(\hat{s}-s)+\sum_{i>1}\lambda_{i}^{k}\langle g,w_{i}\rangle w_{i}$
$E_{s}$ $=\lambda_{1}^{k}\langle s,w_{1}\rangle(\hat{s}-s)+\sum_{i>1}\lambda_{i}^{k}\langle s,w_{i}\rangle w_{i}$
Thus, we have
$x_{k}=\lambda_{1}^{k}\langle s+g,\hat{s}\rangle s+E_{g}+E_{s}$
Thus, we will let $C_{k}=(\lambda_{1}^{k}\langle s+g,\hat{s}\rangle)^{-1}$ and bound the squared norm of the error terms $C_{k}(E_{g}+E_{s})$. We will also use the fact that for any vectors $y,z$, we have $\left\lVert y+z\right\rVert^{2}\leq 2\left\lVert y\right\rVert^{2}+2\left\lVert z\right\rVert^{2}$. Using this, we have
$\left\lVert E_{g}\right\rVert^{2}$ $\leq 2\lambda_{1}^{2k}|\langle\hat{s},g\rangle|^{2}\left\lVert s-\hat{s}\right\rVert^{2}+2\left\lVert\sum_{i>1}\lambda_{i}^{k}\langle g,w_{i}\rangle w_{i}\right\rVert^{2}$
$\leq O(\lambda_{1}^{2k}\delta^{2}/\gamma^{2})|\langle\hat{s},g\rangle|^{2}+2\delta^{2k}\left\lVert\sum_{i>1}\langle g,w_{i}\rangle w_{i}\right\rVert^{2}$
$\leq O(\lambda_{1}^{2k}\delta^{2}/\gamma^{2})|\langle\hat{s},g\rangle|^{2}+2\delta^{2k}\left\lVert g\right\rVert^{2}$
where in the second inequality, we used the fact that Lemma C.1 implies $\left\lVert s-\hat{s}\right\rVert^{2}\leq O(\delta^{2}/\gamma^{2})$. By Lemma A.5, we have $\left\lVert g\right\rVert^{2}=O(\sigma^{2}n\log{n})$ with probability $1-n^{-\Omega(1)}$. Also, $\langle\hat{s},g\rangle$ has distribution $\mathcal{N}(0,1)$ because $\hat{s}$ is a unit vector. Thus with probability $\exp(-\frac{1}{4\sigma^{2}})$, $|\langle\hat{s},g\rangle|\leq 1/2$. If both these events happen, then we have $\left\lVert E_{g}\right\rVert^{2}\leq O(\lambda_{1}^{2k}\delta^{2}/\gamma^{2}+\delta^{2k}\sigma^{2}n\log{n})$. Now, to bound the other error term, we have
$\left\lVert E_{s}\right\rVert^{2}$ $\leq 2\lambda_{1}^{2k}\left\lVert s-\hat{s}\right\rVert^{2}\langle\hat{s},s\rangle^{2}+2\left\lVert\sum_{i>1}\lambda_{2}^{k}\langle w_{i},s\rangle w_{i}\right\rVert^{2}$
$\leq 2\lambda_{1}^{2k}\left\lVert s-\hat{s}\right\rVert^{2}\langle\hat{s},s\rangle^{2}+2\delta^{2k}\left\lVert s\right\rVert^{2}$
$\leq O(\lambda_{1}^{2k}\delta^{2}/\gamma^{2}),$
where the last inequality follows from the fact that $\left\lVert s\right\rVert=1$ and $|\lambda_{1}-\gamma|\leq\delta$, which is a constant factor smaller than $\lambda_{1}$, which means $\delta^{2k}=O(\lambda_{1}^{2k}\delta^{2}/\gamma^{2})$. Finally, we want to lower bound $\langle\hat{s},s\rangle+\langle s,g\rangle$ by a constant so $C_{k}=O(\lambda_{1}^{k})$. By Lemma C.1, we have $\langle\hat{s},s\rangle>1-O(\delta/\gamma)$. We can assume that our constant $C$ is large enough so that $\gamma/\delta<1/4$, which means that $\langle\hat{s},s\rangle\geq 3/4$. Then $\langle\hat{s},s\rangle+\langle s,g\rangle\geq 1/4$ as long as $\langle s,g\rangle\geq 0$ or $|\langle s,g\rangle|\leq 1/2$. This happens with probability at least $\frac{1}{2}+\frac{1}{2}(1-\exp(-\frac{1}{4\sigma^{2}}))=1-\frac{1}{2}\exp(-\frac{1}{4\sigma^{2}})$, and if this does occur, we have
$\left\lVert C_{k}x_{k}-s\right\rVert^{2}$ $\leq 2C_{k}^{2}\left\lVert E_{g}\right\rVert^{2}+2C_{k}^{2}\left\lVert E_{s}\right\rVert^{2}$
$\leq O(\lambda_{1}^{2k}(\lambda_{1}^{-2k}\delta^{2}/\gamma^{2}+\delta^{2k}n\sigma^{2}\log{n}))$
$O(\delta^{2}/\gamma^{2}+(\delta/\lambda_{1})^{2k}n\sigma^{2}\log{n})$
Moreover, since $\lambda_{1}\geq\gamma-\delta$, we have $\delta/\lambda_{1}\leq\delta/(\gamma-\delta)=O(\delta/\gamma)$ since $\gamma$ is a least a constant factor larger than $\delta$, and this gives us our final bound. ∎
With our main proposition established, we can prove our main theorem as follows:
Proof.
(Theorem 4.1) First, we apply Lemma 5.1 to reduce to the one dimensional centered CSBM with parameters $p,q$ and $\sigma^{\prime}=O(\frac{\sigma}{\sqrt{n}\left\lVert\mu-\nu\right\rVert})$. Now let $x=Xw+b$ be our transformed one-dimensional feature vector and let $x_{k}=M^{k}x$ where $M$ is either $\tilde{A}$ or $\hat{A}$. To bound the mean-squared error of our convolved data, we will apply Proposition 6.1. By Proposition 5.2, we can take $\delta=O(\frac{1}{\sqrt{np}})$ if $M=\tilde{A}$ and $\delta=O(\sqrt{\frac{\log{n}}{np}})$ if $M=\hat{A}$. Then, as long as $\gamma\geq C\delta$ for large enough constant $C$, we have a scalar $C_{k}$ and constant $C^{\prime}$ such that
$\left\lVert C_{k}x_{k}-s\right\rVert^{2}\leq O\left(\frac{\delta^{2}}{\gamma^{2}}+\Big{(}\frac{C^{\prime}\delta}{\gamma}\Big{)}^{2k}\frac{\sigma^{2}\log{n}}{\left\lVert\mu-\nu\right\rVert^{2}}\right)$
Now, we take 00 to be the threshold, meaning we put vertex $i$ in the first class if $x_{k}(i)<0$ and put it in the second class otherwise. Note that this partitioning scheme is indifferent to the scaling of the vector $x_{k}$, so we can equivalently apply it to $C_{k}x_{k}$ (note this does not necessitate computing $C_{k}$ directly). Since each $i\in S$ has $s(i)=1/\sqrt{n}$ and each $j\in T$ has $s(j)=-1/\sqrt{n}$, a vertex can only be misclassified if it contributes at least $1/n$ to the total squared distance $\left\lVert C_{k}x_{k}-s\right\rVert^{2}$. Thus, the total number of misclassified vertices is at most $\left\lVert C_{k}x_{k}-s\right\rVert^{2}n$, which gives us the main theorem after substituting the appropriate value of $\delta$. ∎
Appendix D Proofs for Section 7
In this section, we will formally prove Proposition 7.1. Before we begin, we will, as a warm-up, analyze the behavior of 1 convolution for the centered 1-dimensional CSBM. The following proposition is essentially equivalent to theorem 1.2 in Baranwal et al. [2021]. The proof here is not used later on, but gives some intuition about how to analyze exact classification using a matrix-focused framework.
Proposition D.1.
Suppose we are given the 1-dimensional centered CSBM with parameters $n,p,q,\sigma$ such that $p\geq\Omega(\frac{\log^{2}n}{n})$, $\gamma\geq\Omega(\sqrt{\frac{\log{n}}{np}})$, and $\sigma\leq O(\frac{1}{\sqrt{\log{n}}})$. Then with probability $1-n^{-\Omega(1)}$, we have
$\left\lVert s-\frac{1}{\eta}x_{1}\right\rVert_{\infty}\leq O(\sigma\Big{(}\frac{1}{\gamma}\sqrt{\frac{\log{n}}{np}}+\sqrt{\frac{\log{n}}{n}}\Big{)})+\frac{1}{2\sqrt{n}}$
Proof.
By the definition of our convolution matrix, we have
$\frac{1}{\eta}x_{1}=\frac{1}{\eta}(\eta ss^{\top}+R^{\prime})(s+g)=s+\frac{1}{\eta}R^{\prime}s+(ss^{\top}+\frac{1}{\eta}R^{\prime})g$
Thus, we can bound the infinity norm error as:
$\left\lVert s-\frac{1}{\eta}x_{1}\right\rVert_{\infty}=\left\lVert\frac{1}{\eta}R^{\prime}s+(ss^{\top}+\frac{1}{\eta}R^{\prime})g\right\rVert_{\infty}\leq\left\lVert(ss^{\top}+\frac{1}{\eta}R^{\prime})g\right\rVert_{\infty}+\left\lVert\frac{1}{\eta}R^{\prime}s\right\rVert_{\infty}$
Since the infinity norm is the maximum absolute value of all entries, we need to bound terms $|e_{u}^{\top}R^{\prime}s|/\eta$ and $|e_{u}^{\top}(ss^{\top}+R^{\prime}/\eta)g|$ for all $u\in V$. We will start by bounding the Gaussian part of the error. Since $e_{u}^{\top}(ss^{\top}+R^{\prime}/\eta)g\sim\mathcal{N}(0,\sigma^{2}\left\lVert R^{\prime}e_{u}/\eta+\langle s,e_{u}\rangle s\right\rVert^{2})$, we have that with high probability $\max_{i\in n}|e_{u}^{\top}(ss^{\top}+R^{\prime}/\eta)g|\leq O(\sigma\sqrt{\log{n}}\cdot\left\lVert R^{\prime}e_{u}/\eta+\langle s,e_{u}\rangle s\right\rVert)$. Now note that $|\langle s,e_{u}\rangle|=1/\sqrt{n}$ and by Theorem A.4, $\left\lVert R^{\prime}e_{u}\right\rVert\leq\left\lVert R^{\prime}\right\rVert\leq O(\sigma\sqrt{\frac{1}{np}})$. Finally, by Proposition 5.2, we have $\eta=\gamma(1\pm o(1))$. Thus, we have
$|e_{u}^{\top}(ss^{\top}+R^{\prime}/\eta)g|\leq O(\sigma\Big{(}\frac{1}{\gamma}\sqrt{\frac{\log{n}}{np}}+\sqrt{\frac{\log{n}}{n}}\Big{)})$
Now, to bound the error term $|e_{u}^{\top}R^{\prime}s|$, we have $e_{u}^{\top}R^{\prime}s=e_{u}^{\top}(\frac{1}{d}R+d^{\prime}\mathbbm{1}\mathbbm{1}^{\top})s=\frac{1}{d}e_{u}^{\top}Rs$ since $s$ is orthogonal to the all-ones vector. Now, WLOG, suppose $u\in S$. Notice that $e_{u}^{\top}Rs=\frac{1}{\sqrt{n}}(\sum_{i\in S}R_{u,i}-\sum_{i\in T}R_{u,i})$, where both sums are sums over independent shifted Bernoulli random variables with variance at most $p$. Thus, by applying Theorem A.3, we have $|e_{u}^{\top}R^{\prime}s|\leq\frac{1}{\sqrt{n}}\cdot O(\sqrt{np\log{n}})=O(\sqrt{p\log{n}})$. Now, using the fact that $d\geq\frac{1}{2}(p+q)n(1-o(1))$, we have $|\frac{1}{d}e_{u}^{\top}Rs|\leq O(\frac{1}{\sqrt{n}}\cdot\sqrt{\frac{\log{n}}{np}})$. By our assumption, $\gamma\geq C\sqrt{\frac{\log{n}}{np}})$ for large enough constant $C$. Then, we can take $C$ to be large enough so that
$|\frac{1}{\eta}e_{u}^{\top}R^{\prime}s|=O\Big{(}\frac{1}{\gamma(1-o(1))}\sqrt{\frac{\log{n}}{np}}\Big{)}\leq\frac{1}{2\sqrt{n}}$
D.1 Proof of Main Result
In this section, we prove Proposition 7.1, which we restate as follows:
Proposition D.2.
Suppose we are given a 1-dimensional centered 2-block CSBM with parameters $n,p,q,\sigma$ and $k=O(\log{n})$ such that $\gamma=\frac{p-q}{p+q}\geq\Omega(k\sqrt{\frac{\log{n}}{np}})$, $p\geq\frac{\log^{3}{n}}{n}$, and $\sigma\leq O(\frac{1}{\sqrt{\log{n}}})$. Then with probability at least $1-n^{-\Omega(1)}$:
$\left\lVert\frac{1}{\eta^{k}}\tilde{A}^{k}x-s\right\rVert_{\infty}\leq\frac{1}{2\sqrt{n}}+O((\frac{C}{\gamma\sqrt{np}})^{k}\sigma\sqrt{\log{n}})$
Before proving the main proposition, we will use degree-concentration to reduce analyzing the error matrix $R^{\prime}=\tilde{A}-\eta ss^{\top}$ to analysing $\frac{1}{d}R$ instead. This is useful because our analysis crucially uses the fact that our error matrix $R$ has zero-mean Radamacher entries.
Proposition D.3.
Let $R^{\prime}=\tilde{A}-\eta ss^{\top}$ and suppose $k\leq O(\log{n})$. Then with high probability, we have that $\left\lVert R^{\prime k}-\frac{1}{d^{k}}R^{k}\right\rVert\leq O(\frac{C^{k}k\sqrt{\log{n}}}{(\sqrt{pn})^{k}\cdot\sqrt{n}})$ for a constant $C$.
Proof.
Before beginning the proof, we first need to give a tighter degree-concentration bound to show that with high probability, $|d^{\prime}|\leq O(\frac{\sqrt{\log{n}}}{n^{2}\sqrt{p}})$. By applying Theorem A.3 with $t=\Theta(n\sqrt{p\log{n}})$.
$\mathbf{Pr}\left[||E|-\frac{1}{4}(p+q)n^{2}|>t\right]\leq\exp(-\Omega\Big{(}\frac{n^{2}p\log{n}}{(p+q)n^{2}}\Big{)})=n^{-\Omega(1)}$
Since $d=2|E|/n$ and with high probability at, $|E|\geq\frac{1}{4}(p+q)n^{2}-t\geq\frac{1}{4}(p+q)n^{2}(1-o(1))$, we have $d\geq\frac{1}{2}(p+q)n(1-o(1))=\Omega(np)$. Putting these together, we get the bound:
$|d^{\prime}|=\frac{|d-\frac{1}{2}(p+q)n|}{nd}=2\frac{||E|-\frac{1}{4}(p+q)n^{2}|}{n^{2}d}\leq O(\frac{n\sqrt{p\log{n}}}{n^{3}p})\leq O(\frac{\sqrt{\log{n}}}{n^{2}\sqrt{p}})$
Now, recall from Equation 2 that $R^{\prime}=\frac{1}{d}R+d^{\prime}\mathbbm{1}\mathbbm{1}^{\top}$. By Proposition 5.2, we have that with high probability, $\frac{1}{d}\left\lVert R\right\rVert\leq\sqrt{\frac{C}{np}}$ for some constant $C$. We will let $\delta:=\sqrt{\frac{C}{np}}$ be this upper bound. Now consider the matrix $(\frac{1}{\delta}R^{\prime})^{k}$. We will let $M_{0}=\frac{1}{d\delta}R$ and $M_{1}=\frac{d^{\prime}}{\delta}\mathbbm{1}\mathbbm{1}^{\top}$. Then, we have $\left\lVert M_{0}\right\rVert\leq 1$, and $\left\lVert M_{1}\right\rVert\leq\frac{d^{\prime}n}{\delta}$. Since $|d^{\prime}|\leq\frac{\sqrt{\log{n}}}{n^{2}\sqrt{p}}$ with high probability, we will assume this event occurs. Thus, we have $\left\lVert M_{1}\right\rVert\leq O(\frac{n\sqrt{np\log{n}}}{n^{2}\sqrt{p}})\leq O(\sqrt{\frac{\log{n}}{n}})$. Finally, we have $(\frac{1}{\delta}R^{\prime})^{k}=(M_{0}+M_{1})^{k}$. We are going to expand this out, and so let us introduce some notation. Let $\binom{[k]}{\ell}=\{i:=(i_{1},i_{2},...i_{k})\in\{0,1\}^{k}:i_{1}+i_{2}+...i_{k}=\ell\}$. That is, it’s the set of length $k$ binary sequences with exactly $\ell$ terms equaled to $1$. Now, we have
$(M_{0}+M_{1})^{k}-M_{0}^{k}$ $=\sum_{\ell=0}^{k}\sum_{i\in\binom{[k]}{\ell}}\prod_{j=1}^{k}M_{i_{j}}-M_{0}^{k}$
$=\sum_{\ell=1}^{k}\sum_{i\in\binom{[k]}{\ell}}\prod_{j=1}^{k}M_{i_{j}}$
For any matrices $M,N$, their spectral norms satisfy $\left\lVert MN\right\rVert\leq\left\lVert M\right\rVert\left\lVert N\right\rVert$. Thus, for each $i\in\binom{[k]}{\ell}$, we have $\left\lVert\prod_{j=1}^{k}M_{i_{j}}\right\rVert\leq\left\lVert M_{1}\right\rVert^{\ell}$ since $\left\lVert M_{0}\right\rVert\leq 1$. Thus, we have
$\left\lVert(M_{0}+M_{1})^{k}-M_{0}^{k}\right\rVert$ $\leq\sum_{\ell=1}^{k}\binom{k}{\ell}\left\lVert M_{1}\right\rVert^{\ell}$
$=(1+\left\lVert M_{1}\right\rVert)^{k}-1$
$\leq e^{k\left\lVert M_{1}\right\rVert}-1$
$\leq 2k\left\lVert M_{1}\right\rVert$
where the second last inequality follows from Lemma A.6 because $k\left\lVert M_{1}\right\rVert\leq k\sqrt{\frac{\log{n}}{n}}\leq 1$ by our assumptions on $k$. Finally, we have
$\left\lVert R^{\prime k}-\frac{1}{d^{k}}R^{k}\right\rVert$ $=\delta^{k}\left\lVert(M_{0}+M_{1})^{k}-M_{0}\right\rVert$
$\leq 2\delta^{k}k\left\lVert M_{1}\right\rVert$
$\leq O(\frac{kC^{k}\sqrt{\log{n}}}{(\sqrt{pn})^{k}\cdot\sqrt{n}})$
For a constant $C$. ∎
As a corollary, we can show the following:
Corollary D.4.
Let $x$ and $y$ be unit vectors. Then $|x^{\top}R^{\prime k}y|\leq|x^{\top}(\frac{1}{d^{k}}R^{k})y|+O(\frac{C^{k}k\sqrt{\log{n}}}{(\sqrt{pn})^{k}\cdot\sqrt{n}}))$ for some constant $C$.
Proof.
This follows from the basic properties of the spectral norm
$|x^{\top}R^{\prime k}y|$ $=|x^{\top}(\frac{1}{d^{k}}R^{k}+R^{\prime k}-\frac{1}{d^{k}}R^{k})y|$
$\leq|x^{\top}(\frac{1}{d^{k}}R^{k})y|+\left\lVert x\right\rVert\left\lVert y\right\rVert\left\lVert R^{\prime k}-\frac{1}{d^{k}}R^{k}\right\rVert$
[…]
Corollary D.4 is key to proving our main technical lemma, the proof of which we will defer to the next section.
Lemma D.5.
Given $p\geq\frac{\log^{3}{n}}{n}$, we have that with probability $1-n^{-\Omega(1)}$, for all $u\in V$ and all $k\in\{1,2,...O(\log{n})\}$
$|e_{u}^{\top}R^{\prime k}s|\leq\frac{1}{\sqrt{n}}\Big{(}C\sqrt{\frac{\log{n}}{pn}}\Big{)}^{k}$
for some constant $C$
Now, we are ready to prove our main result, Proposition 7.1. The proof is similar to that of Proposition D.3 but requires the use of Lemma D.5 to obtain a sharper bound on the error terms than by simply applying the spectral norm bound as we did in the degree-concentration proof.
Proof.
(Proposition 7.1) Just like in the $k=1$ case, we express $\tilde{A}$ as $\eta ss^{\top}+R^{\prime}$ and decompose our convolution vector into the following terms:
$\frac{1}{\eta^{k}}\tilde{A}^{k}x$ $=\frac{1}{\eta^{k}}\tilde{A}^{k}(s+g)$
$=s+(\frac{1}{\eta^{k}}\tilde{A}^{k}-ss^{\top})s+\frac{1}{\eta^{k}}\tilde{A}^{k}g$
In order to bound $\left\lVert s-\frac{1}{\eta^{k}}\tilde{A}^{k}x\right\rVert_{\infty}$, we will bound, for all $u\in V$, the error terms $|e_{u}^{\top}(\frac{1}{\eta^{k}}\tilde{A}^{k}-ss^{\top})s|$ and $|e_{u}^{\top}\frac{1}{\eta^{k}}\tilde{A}^{k}g|$. Similar to in the proof of Proposition D.3, we start by expanding out $\frac{1}{\eta^{k}}\tilde{A}^{k}$. Let $Y_{0}=ss^{\top}$ and $Y_{1}=R^{\prime}$. Then we have
$\frac{1}{\eta^{k}}\tilde{A}^{k}$ $=\frac{1}{\eta^{k}}(\eta Y_{0}+Y_{1})^{k}$
$=\frac{1}{\eta^{k}}\sum_{\ell=0}^{k}\eta^{k-\ell}\sum_{i\in\binom{[k]}{\ell}}\prod_{j=1}^{k}Y_{i_{j}}$
$=\sum_{\ell=0}^{k}\eta^{-\ell}\sum_{i\in\binom{[k]}{\ell}}\prod_{j=1}^{k}Y_{i_{j}}$
Note that $(ss^{\top})^{2}=ss^{\top}$, which means that the first term in our summation (i.e. the $\ell=0$ terms) is simply $ss^{\top}$. Now we start by bounding the error terms involving $s$.
$|e_{u}^{\top}(\frac{1}{\eta^{k}}\tilde{A}^{k}-ss^{\top})s|$ $=|\sum_{\ell=1}^{k}\eta^{-\ell}\sum_{i\in\binom{[k]}{\ell}}e_{u}^{\top}\prod_{j=1}^{k}Y_{i_{j}}s|$
$\leq\sum_{\ell=1}^{k}\eta^{-\ell}\sum_{i\in\binom{[k]}{\ell}}|e_{u}^{\top}\prod_{j=1}^{k}Y_{i_{j}}s|$
For each $i\in\binom{[k]}{\ell}$, we have
$|e_{u}^{\top}\prod_{j=1}^{k}Y_{i_{j}}s|$ $=|e_{u}^{\top}R^{\prime a_{1}}s\cdot s^{\top}R^{\prime a_{2}}s\cdot...s^{\top}R^{\prime a_{L}}s|$
where $a_{1},...a_{L}$ are non-negative integers satisfying $a_{1}+...a_{L}=\ell$. This is because the product $\prod_{j=1}^{k}Y_{i_{j}}$ has exactly $\ell$ terms of the form $R^{\prime}$ and $ss^{\top}$ for the rest of the terms. By Proposition 5.2, the matrix $R^{\prime}$ has spectral norm at most $\frac{C_{1}}{\sqrt{np}}$ for some constant $C_{1}$ with high probability. For each term of the form $|s^{\top}R^{\prime a_{j}}s|$ for $j>1$, we bound it by $|s^{\top}R^{\prime a_{j}}s|\leq\left\lVert R^{\prime}\right\rVert^{a_{j}}\leq\Big{(}\frac{C_{1}}{\sqrt{np}}\Big{)}^{a_{j}}$. For the first term $|e_{u}^{\top}R^{\prime a_{1}}s|$, we apply Lemma D.5 and give $|e_{u}^{\top}R^{\prime a_{1}}s|\leq\frac{1}{\sqrt{n}}\Big{(}C_{2}\sqrt{\frac{\log{n}}{np}}\Big{)}^{a_{1}}$ for some constant $C_{2}$. Thus, we have:
$|e_{u}^{\top}\prod_{j=1}^{k}Y_{i_{j}}s|\leq\frac{1}{\sqrt{n}}\Big{(}\frac{\max(C_{1},C_{2})}{\sqrt{np}}\Big{)}^{a_{1}+...a_{L}}\sqrt{\log{n}}^{a_{1}}\leq\frac{1}{\sqrt{n}}\Big{(}C\sqrt{\frac{\log{n}}{np}}\Big{)}^{\ell}$
where $C$ is a large enough constant such that $C\geq\max(C_{1},C_{2})$. Now, we let $\rho:=\frac{C}{\eta}\sqrt{\frac{\log{n}}{np}}$. Assuming that $\gamma\geq 9Ck\sqrt{\frac{\log{n}}{np}}$ and with high probability, $\eta\in\gamma(1\pm o(1))$, we have $\rho\leq\frac{1}{8k}$ with high probability. Thus, we have
$|e_{u}^{\top}(\frac{1}{\eta^{k}}\tilde{A}^{k}-ss^{\top})s|$ $\leq\sum_{\ell=1}^{k}\eta^{-\ell}\sum_{i\in\binom{[k]}{\ell}}|e_{u}^{\top}\prod_{j=1}^{k}Y_{i_{j}}s|$
$\leq\frac{1}{\sqrt{n}}\sum_{\ell=1}^{k}\binom{k}{\ell}\eta^{-\ell}\Big{(}C\sqrt{\frac{\log{n}}{np}}\Big{)}^{\ell}$
$=\frac{1}{\sqrt{n}}\sum_{\ell=1}^{k}\binom{k}{\ell}\rho^{\ell}$
$=\frac{1}{\sqrt{n}}((1+\rho)^{k}-1)$
[…]
Now, we bound the Gaussian part of the error:
$e_{u}^{\top}\tilde{A}g=\sum_{\ell=0}^{k-1}\eta^{-\ell}\sum_{i\in\binom{[k]}{\ell}}e_{u}^{\top}\prod_{j=1}^{k}Y_{i_{j}}g+\eta^{-k}e_{u}^{\top}R^{\prime k}g$
For each $i\in\binom{[k]}{\ell}$ where $\ell<k$, we have
$|e_{u}^{\top}\prod_{j=1}^{k}Y_{i_{j}}g|=|e_{u}^{\top}R^{\prime a_{1}}s\cdot s^{\top}R^{\prime a_{2}}s\cdot...s^{\top}R^{\prime a_{L}}g|$
where $a_{1}+a_{2}+...a_{L}=\ell$. This is because when $\ell<k$, there is at least one $Y_{i_{j}}$ term that is equalled to $ss^{\top}$, which means in the product, the leftmost term must be $e_{u}^{\top}R^{\prime a_{1}}s$ for some $a_{1}$ and the right most term must be $s^{\top}R^{\prime a_{L}}g$ for some $a_{L}$. Once again, we use the fact that $|s^{\top}R^{\prime a_{j}}s|\leq\left\lVert R^{\prime}\right\rVert^{a_{j}}\leq(\frac{C_{1}}{\sqrt{np}})^{a_{j}}$ for all $j>1$, and by Lemma D.5, we have $|e_{u}^{\top}R^{\prime a_{1}}s|\leq(C_{2}\sqrt{\frac{\log{n}}{np}})^{a_{1}}$ where $C_{1},C_{2}$ are absolute constants. To bound the last term, we have that $s^{\top}R^{\prime a}g\sim\mathcal{N}(0,\sigma^{2}\left\lVert R^{\prime a}s\right\rVert^{2})$ for all $a>0$. Thus, with high probability, we have that $|s^{\top}R^{\prime a_{L}}g|\leq O(\sigma\left\lVert R^{\prime a_{L}}s\right\rVert\sqrt{\log{n}})\leq O(\sigma(\frac{C_{1}}{\sqrt{np}})^{a_{L}}\sqrt{\log{n}})$. Thus, we can apply the same bounds as we did for the error term involving $s$ but now with an extra $\sigma\sqrt{\log{n}}$ factor:
$|e_{u}^{\top}\prod_{j=1}^{k}Y_{i_{j}}g|$ $=|e_{u}^{\top}R^{\prime a_{1}}s\cdot s^{\top}R^{\prime a_{2}}s\cdot...s^{\top}R^{\prime a_{L}}g|$
$\leq O(\frac{\sigma\sqrt{\log{n}}}{\sqrt{n}}\cdot\sum_{\ell=0}^{k-1}\binom{k}{\ell}\rho^{\ell})$
$\leq O(\frac{\sigma\sqrt{\log{n}}}{\sqrt{n}}\cdot(1+\rho)^{k})$
[…]
$\leq O(\frac{1}{4\sqrt{n}})$
Where the last inequality follows assuming $\sigma\leq\frac{1}{C^{\prime}\sqrt{\log{n}}}$ for some large enough constant $C^{\prime}$. Finally, to bound the $k^{th}$ order term, we have that with high probability, $\eta^{-k}|e_{u}^{\top}R^{\prime k}g|\leq(\frac{C_{1}}{\eta\sqrt{np}})^{k}\sigma\sqrt{\log{n}}$. Putting everything together, and using the fact that $\eta\geq\gamma(1-o(1))$ with high probability, we have that for all $u\in V$,
$|e_{u}^{\top}s-\frac{1}{\eta^{k}}\tilde{A}^{k}x|\leq\frac{1}{\eta^{k}}(|e_{u}^{\top}(\tilde{A}^{k}-ss^{\top})s|+|e_{u}^{\top}\tilde{A}^{k}g|)\leq\frac{1}{2\sqrt{n}}+\Big{(}\frac{C}{\gamma\sqrt{np}}\Big{)}^{k}\sigma\sqrt{\log{n}})$
for some absolute constant $C$
Given Proposition 7.1, we can derive Theorem 4.2 by using the standard reduction in Lemma 5.1.
Proof.
(Theorem 4.2) By applying Lemma 5.1, we see that given $m$-dimensional features with feature matrix $X$, we can transform it to a centered 1-dimensional feature vector $x=Xw+b=s+g$ where $g\sim\mathcal{N}(0,\sigma^{\prime 2}I)$ and $\sigma^{\prime}=\frac{4\sigma}{\sqrt{n}\left\lVert\nu-\mu\right\rVert}$. Thus, we have $\left\lVert\nu-\mu\right\rVert=\frac{4\sigma}{\sigma^{\prime}\sqrt{n}}$. By Proposition 7.1, our 1-dimensional features become linearly separable as long as $\Big{(}\frac{C}{\gamma\sqrt{np}}\Big{)}^{k}\sigma^{\prime}\sqrt{\log{n}}<\frac{1}{2\sqrt{n}}$ for some absolute constant $C$. Given our expression of $\sigma^{\prime}$ in terms of the mean distance, this is equivalent to
$\left\lVert\nu-\mu\right\rVert\geq\sigma\Big{(}\frac{C}{\gamma\sqrt{np}}\Big{)}^{k}\sqrt{\log{n}}$
In Proposition 7.1, we also needed to assumed that $\sigma^{\prime}\leq O(\frac{1}{\sqrt{\log{n}}})$, which implies that
$\left\lVert\nu-\mu\right\rVert\geq\Omega(\sigma\sqrt{\frac{\log{n}}{n}})$
D.2 Message Passing Error Bound
In this section, we will prove our main technical lemma, Lemma D.5. By Corollary D.4, it suffices to control the term $|e_{u}^{\top}R^{k}s|$ in order to control $|e_{u}^{\top}R^{\prime k}s|$. Since this is the sum over many dependent random variables, we cannot easily compute its moment generating function. Instead, we will compute the moments directly. In particular, we will apply Markov’s inequality using the $2t^{th}$ moment of this random variable $\mathbb{E}[(e_{u}^{\top}R^{k}s)^{2t}]$ for an appropriately chosen $t$. We now state our main result for this section as follows:
Proposition D.6.
Suppose $R$ is an $n\times n$ symmetric random matrix where for each $1\leq i\leq j$, $R_{i,j}=1-p_{i,j}$ with probability $p_{i,j}$ and $-p_{i,j}$ with probability $1-p_{i,j}$ and are independent. Let $p=\max_{i,j}p_{i,j}$, and suppose $p\geq\Omega(\frac{\log^{3}{n}}{n})$. Then we have
$\mathbf{Pr}\left[|e_{u}^{\top}R^{k}s|>\frac{1}{\sqrt{n}}(C\sqrt{np\log{n}})^{k}\right]\leq\exp(-\Omega(C\log{n}))$
Proof.
The term $e_{u}^{\top}R^{k}s$ can be written as the sum over all walks of length $k$ originating from $u$. Let $\mathcal{W}_{u,k}$ denote the set of all walks of length $k$ originating at $u$. For brevity, we will use $\mathcal{W}$ to denote $\mathcal{W}_{u,k}$. For each $w\in\mathcal{W}$, let $w(j)$ be the $j^{th}$ vertex in the walk. Note that $w(0)=u$ always. Then we have
$e_{u}^{\top}R^{k}s$ $=\sum_{w\in\mathcal{W}}\prod_{j=1}^{k}R_{w(j-1),w(j)}s(w(k))$
Thus, we have
$\mathbb{E}[(e_{u}^{\top}R^{k}s)^{2t}]$ $=\sum_{w_{1},w_{2}...w_{2t}\in\mathcal{W}}\mathbb{E}[\prod_{i=1}^{2t}\prod_{j=1}^{k}R_{w_{i}(j-1),w_{i}(j)}s(w_{i}(k))]$
$\leq\sum_{w_{1},w_{2}...w_{2t}\in\mathcal{W}}|\mathbb{E}[\prod_{i=1}^{2t}\prod_{j=1}^{k}R_{w_{i}(j-1),w_{i}(j)}]|\cdot\prod_{i=1}^{2t}|s(w_{i}(k))|$
$=\frac{1}{n^{t}}\sum_{w_{1},w_{2}...w_{2t}\in\mathcal{W}}|\mathbb{E}[\prod_{i=1}^{2t}\prod_{j=1}^{k}R_{w_{i}(j-1),w_{i}(j)}]|$
Where the last equality follows from the fact that $|s(v)|=1/\sqrt{n}$ for all $v\in V$. Now, for each $\vec{w}=(w_{1},...w_{2t})\in\mathcal{W}^{2t}$ and edge $h\in n\times n$, let #$\#_{\vec{w}}(h)$ be the number of times the edge $h$ occurs in the graph $w_{1}\cup w_{2}\cup,...w_{2t}$. In other words, for $h=\{a,b\}$, #$\#_{\vec{w}}(h)$ is the number of times the variable $R_{a,b}$ occurs in the product $\prod_{i=1}^{2t}\prod_{j=1}^{k}R_{w_{i}(j-1),w_{i}(j)}$. We will also let $|\vec{w}|$ denote the number of unique edges in the graph formed by the union of these $2t$ walks. For example, if $\vec{w}$ consist of the walks $(1,2,3)$, and $(1,2,4)$, then we have #$\#_{\vec{w}}(1,2)=2$, #$\#_{\vec{w}}(2,3)=1$, #$\#_{\vec{w}}(2,4)=1$, and $|\vec{w}|=3$. Using this notation, we have
$\sum_{w_{1},w_{2}...w_{2t}\in\mathcal{W}}|\mathbb{E}[\prod_{i=1}^{2t}\prod_{j=1}^{k}R_{w_{i}(j-1),w_{i}(j)}]|$ $=\sum_{\vec{w}\in\mathcal{W}^{2t}}\prod_{h\in[n]\times[n]}|\mathbb{E}[R_{h}^{\#_{\vec{w}}(h)}]|$
Now we note that each $R_{h}$ has $\mathbb{E}[R_{h}]=0$ and for all $k\geq 2$, we have
$|\mathbb{E}[R_{h}^{k}]|\leq\mathbb{E}[|R_{h}^{k}|]=p_{h}(1-p_{h})^{k}+p_{h}^{k}(1-p_{h})=p_{h}((1-p_{h})^{k}+p_{h}^{k-1}(1-p_{h}))\leq p_{h}\leq p$
Thus, the term $\prod_{h\in[n]\times[n]}|\mathbb{E}[R_{h}^{\#_{\vec{w}}(h)}]|$ is only nonzero if in the union of the edges in the walks $w_{1},...w_{2t}$, each edge is counted at least twice, and if it is non-zero, then it is the product of at most $|\vec{w}|$ terms of value at most $p$. We now let $\mathcal{W}_{pair}^{2t}$ be the set of such walks, which we will denote as “valid walks”. As an example, when $t=2$ and $k=2$, the walks $\{(1,2,3),(1,2,3)\}$ would be valid but $\{(1,2,3),(1,2,4)\}$ would not be valid. Now we have
$\sum_{\vec{w}\in\mathcal{W}^{2t}}\prod_{h\in[n]\times[n]}|\mathbb{E}[R_{h}^{\#_{\vec{w}}(h)}]|$ $\leq\sum_{w\in\mathcal{W}_{pair}^{2t}}p^{|\vec{w}|}$
$=\sum_{\ell=1}^{tk}p^{\ell}|\{\vec{w}\in\mathcal{W}_{pair}^{2t},\;|\vec{w}|=\ell\}|$
Now, what we have left is a counting problem. We need to count the number of valid sets of walks whose union has exactly $\ell$ distinct edges. We note that $\ell$ can be at most $tk$ because otherwise, there must be an edge that is counted at most once, making the walks invalid. Since it is difficult to count this quantity exactly, we will just upper-bound it as follows:
Proposition D.7.
Suppose $p\geq\frac{\log^{3}{n}}{n}$ and $t\leq\frac{\log{n}}{2k}$. Then we have:
|$|\{\vec{w}\in\mathcal{W}_{pair}^{2t},\;|\vec{w}|=\ell\}|\leq\binom{2tk}{2\ell}(2\ell-1)!!\cdot\ell^{2kt-2\ell}(n)^{\ell}$
The notation of $n!!$! denotes $n\cdot n-2\cdot n-4...1$. Given this upper bound on the number of valid sets of walks, we are ready to give our final bound. We will let $t=\frac{C_{1}\log{n}}{2k}$ for large enough constant $C_{1}$. Note that the only constraint on $t$ is that it is a positive integer, and this is feasible as long as $k=O(\log{n})$. Since $p\geq\frac{\log^{3}n}{n}$, we have $np\geq(tk/C_{1})^{3}$. We will make the substitution $b=tk-\ell$. Then we have
$\sum_{\ell=1}^{tk}p^{\ell}|\{\vec{w}\in\mathcal{W}_{pair}^{2t},\;|\vec{w}|=\ell\}|$ $\leq\sum_{b=0}^{tk-1}(np)^{tk-b}(tk-b)^{2b}(2tk-2b-1)!!\binom{2tk}{2b}$
$\leq\sum_{b=0}^{tk-1}(np)^{tk-b}(2tk)^{tk-b}(2tk)^{2b}\frac{(2tk)^{2b}e^{2b}}{(2b)^{2b}}$
$\leq(np)^{tk}(2tk)^{tk}\sum_{b=0}^{tk-1}\frac{(2tk)^{3b}e^{2b}}{(2b)!(np)^{b}}$
$\leq(np)^{tk}(2tk)^{tk}\sum_{b\geq 0}\frac{(C_{1}e^{2})^{b}}{(2b)!}$
$\leq O((np)^{tk}(2tk)^{tk})$
In the second line, we used the fact that $(2tk-2b-1)!!$! is the product of $tk-b$ terms of value at most $2tk$ as well as the standard upper bound of $\binom{2tk}{2b}\leq\frac{(2tk)^{2b}e^{2b}}{(2b)^{2b}}$, the second last bound follows from our assumption on $p$, and the final bound follows from the fact that the sum converges. Putting all of this together, we have that for $t=\frac{C_{1}\log{n}}{2k}$ and $p\geq\frac{\log^{3}n}{n}$,
$\mathbb{E}[(e_{u}^{\top}R^{k}s)^{2t}]\leq O(\frac{(2tknp)^{tk}}{n^{t}})$
Now, we can apply Markov’s inequality on the $t^{th}$ moment with $t=\frac{C_{1}\log{n}}{2k}$
$\mathbf{Pr}\left[[\right]|e_{u}^{\top}R^{k}s|>\frac{1}{\sqrt{n}}(C\sqrt{np\log{n}})^{k}]$ $\leq\frac{n^{t}\mathbb{E}[(e_{u}^{\top}R^{k}s)^{2t}]}{C^{2tk}(np\log{n})^{tk}}$
$\leq O(\frac{(np\log{n})^{tk}C_{1}^{tk}}{C^{2tk}(np\log{n})^{tk}})$
$\leq O(\frac{\sqrt{C_{1}}^{\log{n}}}{C^{\log{n}}})$
$=\exp(-\Omega(\log{n}))$
for a sufficiently large constant $C$
Now, we combine our results from degree-concentration and the message passing error bound to prove the main result of this section, Lemma D.5.
Proof.
(Lemma D.5) First, note that for $k=0$, the bound clearly holds, as $|e_{u}^{\top}s|=\frac{1}{\sqrt{n}}$. Now for general $k$, we start by applying Corollary D.4 to obtain
$|e_{u}^{\top}R^{\prime k}s|$ $\leq\frac{1}{d^{k}}|e_{u}^{\top}R^{k}s|+O(\frac{kC_{1}^{k}\sqrt{\log{n}}}{\sqrt{np}^{k}\cdot\sqrt{n}})$
For some constant $C_{1}$. Now by Proposition D.6, we have that with high probability $e_{u}^{\top}R^{k}s\leq\frac{1}{\sqrt{n}}(C_{2}\sqrt{np\log{n}})^{k}$ for some constant $C_{2}$ and $d\geq\frac{1}{2}(p+q)n(1-o(1))$. Thus, we have $\frac{1}{d^{k}}|e_{u}^{\top}R^{k}s|\leq\frac{1}{\sqrt{n}}\Big{(}C_{2}\sqrt{\frac{\log{n}}{pn}}\Big{)}^{k}$. Note that $\sqrt{\log{n}}^{k}$ is bigger than $k\sqrt{\log{n}}$ for large enough $n$. Thus, the $\frac{1}{d^{k}}|e_{u}^{\top}R^{k}s|$ term clearly dominates so we can take $C$ to be around $\max(C_{1},C_{2})$ to obtain our upper bound. Finally, by applying union bound, we can ensure that with high probability, bound holds for all $u\in V$. ∎
D.2.1 Proof of proposition Proposition D.7
In this section, we complete our proof of the main message passing error bound by proving the upper bound on the number of valid sets of walks of length $k$.
Proof.
(Proposition D.7) The idea to bound the quantity |$|\{\vec{w}\in\mathcal{W}_{pair}^{2t},\;|\vec{w}|=\ell\}|$ is to first count all the ways to partition the edges of the $2t$ walks into exactly $\ell$ groups of size at least $2$. Then, for each partition, we count the number of ways to assign the vertices of the walks in a way such that all edges of the same group are assigned to the same pair of vertices. In the grouping stage, we first pick $2\ell$ walk edges and pair them with each other, assigning a new group to each pair. Then for the rest of the $2tk-2\ell$ edges, we assign each of them to one of the $\ell$ groups. This ensures that each group has at least two edges. The number of ways to pick $2\ell$ edges is $\binom{2kt}{2\ell}$. The number of ways to pair these edges is $(2\ell-1)!!=(2\ell-1)(2\ell-3)...\cdot 1$. Finally, the number of way to assign the rest of the edges to one of the groups is $\ell^{(2kt-2\ell)}$. Thus, the total number of valid partitions of the edges is at most $\binom{2tk}{2\ell}(2\ell-1)!!\cdot\ell^{2kt-2\ell}$. Note that this upper bound is tight when $\ell=tk$ but becomes less tight as $\ell$ becomes smaller.
Given a partition of the edges, we now count the number of ways to assign the vertex variables $w_{i}(j)$. We will give a simple assignment algorithm such that every valid assignment is a possible output of the algorithm. Note that not every output of the algorithm is necessarily valid, so we are actually over-counting.
Vertex Assignment Procedure
  • Pick an arbitrary partitioning of the edges into $\ell$ groups without singletons
  • For $i=1$ to $2t$:
    • For $j=1$ to $k$
        If the edge $(w_{i}(j-1),w_{i}(j))$ is not part of a group which contains an already assigned edge, then pick a new vertex for $w_{i}(j)$.
        Else, let $(w_{i^{\prime}}(j^{\prime}-1),w_{i^{\prime}}(j^{\prime}))$ be an edge in the same group as $(w_{i}(j-1),w_{i}(j))$ such that both $w_{i^{\prime}}(j^{\prime}-1)$ and $w_{i^{\prime}}(j^{\prime})$ are already assigned. Then assign $w_{i}(j)$ such that $\{w_{i}(j-1),w_{i}(j)\}=\{w_{i^{\prime}}(j^{\prime}-1),w_{i^{\prime}}(j^{\prime})\}$. If no such assignment is possible, then the procedure fails.
To see that every possible valid set of walks can be outputted by this procedure, if we are given some $\vec{w}=(w_{1},...w_{2t})\in\mathcal{W}_{pair}^{2t}$, we can take the partitioning of the edges in the first step to be to be the partition induced by the union of the edges of the $2t$ walks. Then, we can simply follow the vertices of the walks in the same order as in our assignment procedure and see that each time we encounter a new edge corresponds to the first case of the inner for-loop and each time we traverse an already traversed edge corresponds to the second case in the inner for-loop where the procedure does not fail. Thus, the number of valid sets of walks with $\ell$ distinct edges is indeed upper bounded by the total number of possible outputs to our vertex assignment procedure.
To bound the number of possible outcomes in our procedure, we see that in first case of the inner for-loop, there are at most $n$ possible assignments for the vertex $w_{i}(j)$. In the second step, there is only one possible assignment if the procedure doesn’t fail. Since we encounter the first case of the for-loop at most $\ell$ times, the total number of vertex assignments given a fixed partitioning of edges is at most $n^{\ell}$. Finally, this gives us the desired bound:
|$|\{\vec{w}\in\mathcal{W}_{pair}^{2t},\;|\vec{w}|=\ell\}|\leq\binom{2tk}{2\ell}(2\ell-1)!!\cdot\ell^{2kt-2\ell}(n)^{\ell}$